### v2

# Generative modeling of conditional probability distributions on the level-sets of collective variables

Abstract
Given a probability distribution $\mu$ in $\mathbb{R}^{d}$ represented by data, we study in this paper the generative modeling of its conditional probability distributions on the level-sets of a collective variable $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$. We propose a general and efficient learning approach that is able to learn generative models on different level-sets of $\xi$ simultaneously. To improve the learning quality on level-sets in low-probability regions, we also propose a strategy for data enrichment by utilizing data from enhanced sampling techniques. We demonstrate the effectiveness of our proposed learning approach through concrete numerical examples. The proposed approach is potentially useful for the generative modeling of molecular systems in biophysics, for instance.
Keywords— collective variables, level-sets, conditional probability distributions, generative modeling, flow-matching
1 Introduction
High-dimensional stochastic processes are ubiquitous in broad research areas such as biophysics, materials science, and climate modeling. Understanding the complex dynamical behavior of such processes often requires projecting the system down to a (much) lower-dimensional space via a few features, i.e. collective variables (CVs) or reaction coordinates, of the system. In molecular dynamics, for instance, various enhanced sampling techniques have been developed to overcome the sampling difficulty due to the system’s metastability and to construct the free energy surface in the space of CVs [15]. A number of analytical and computational approaches have also been proposed for constructing simpler, lower-dimensional surrogate models associated to a chosen set of CVs of the original high-dimensional system. One notable approach is the effective dynamics constructed using conditional expectations, where the coefficients (i.e. drift and diffusion) in the surrogate model are computed by taking averages with respect to the so-called conditional probability distribution on the corresponding level-set of the CVs [21]. This approach has attracted considerable research interest in the mathematical community [50, 51, 22, 29, 23, 9, 52], contributing to a deeper understanding of the quality of the effective dynamics in approximating the original high-dimensional processes.
A related research topic in molecular dynamics is sampling on the level-sets of CVs [5, 13, 26], which is useful in both of the aforementioned tasks, i.e. free energy computation and computing the coefficients in the effective dynamics. Metropolis-Hastings Monte Carlo methods have been developed for unbiased sampling on submanifolds defined by the level-sets of CVs [48, 27, 28]. Given a (Boltzmann–Gibbs) distribution in the full space, sampling methods for its conditional distribution on the level-sets of CVs have been developed by constructing ergodic diffusion processes on the level-sets [49, 42]. More broadly, methods for sampling under constraints have also been studied in computational statistics [46, 35].
In recent years, generative modeling approaches, in particular diffusion models [43, 17, 44, 45] and flow-matching models [31, 32, 1], have achieved significant success in a wide range of real-world applications. Motivated by these achievements, there has been growing research interest in extending generative modeling approaches for Euclidean data to manifold data [36, 41, 2, 8, 18, 19, 4, 34, 33]. However, to the best of our knowledge, most of these existing approaches rely on substantial geometric information such as geodesics [8], heat kernel, or eigenfunctions [4], which consequently limits their applicability to manifolds where such information is explicitly known. Only the methods proposed in [34, 33] can be applied to manifolds that are implicitly defined as a level-set of a certain function. All these approaches are developed for generative modeling on a single and fixed manifold.
In this paper, we study the generative modeling of conditional probability distributions on the level-sets of CVs and propose a versatile framework for this task. Our approach builds upon the flow-matching model [31, 32, 1], in which a neural ordinary differential equation (ODE) is learned to transport samples from a base Gaussian distribution to the target conditional distribution on each level-set of a CV map $\xi$. Our method does not require explicit geometric information of the underlying manifolds and therefore is broadly applicable to a wide range of manifolds implicitly defined by CVs. Moreover, by varying the target CV value, the method can generate samples across different level-sets of $\xi$, rather than being restricted to a single level-set. To further improve learning in regions of low probability density, we introduce a data enrichment strategy that leverages data obtained from enhanced sampling techniques such as adaptive biasing force (ABF) simulations. The approach is validated on a hierarchy of systems of increasing complexity, from synthetic low-dimensional examples to high-dimensional molecular datasets. The main contributions of this paper are summarized below:
  1. (i)
    a general and efficient generative modeling framework capable of learning conditional distributions on families of manifolds implicitly defined by CVs;
  2. (ii)
    a data enrichment strategy that utilizes enhanced sampling data to improve accuracy in low-density regions; and
  3. (iii)
    comprehensive numerical validations that demonstrate the flexibility, robustness, and practical applicability of the method to molecular systems.
The remainder of the paper is organized as follows. In Section 2, we present the theoretical background and introduce the relevant notation. Section 3 details the proposed generative modeling methodology. In Section 4, we apply our approach to a series of representative datasets to illustrate its performance and generality. Finally, Section 5 concludes the paper and outlines possible directions for future research.
2 Background and notation
The aim of this section is to introduce the probability distributions of interest in this work, in particular the conditional distribution on the level-sets of a given smooth map, which will play a central role in the following.
Consider a target distribution $\mu$ over $\mathbb{R}^{d}$ with a positive probability density $\rho$ with respect to the Lebesgue measure, namely $\mu(dx)=\rho(x)\,dx$. Assume that a $C^{2}$-smooth map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is given, where $1\leq k<d$. We are particularly interested in applications in statistical physics, where $\mu$ arises as the invariant distribution of certain ergodic diffusion processes $x_{t}$ in $\mathbb{R}^{d}$ and $\xi$ corresponds to the CV of the system. In this case, we often refer to the Boltzmann–Gibbs distribution, whose density with respect to the Lebesgue measure is
$\rho(x)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)},\quad x\in\mathbb{R}^{d}\,$, (1)
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that grows sufficiently fast to infinity as $|x|\rightarrow+\infty$, $\beta$ is a positive constant whose inverse is proportional to the system’s temperature, and $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,dx$ is the normalizing constant. We introduce a few quantities related to the measure $\mu$ and the map $\xi$. Further discussions can be found in [25, Section 3.2.1] and [50].
Given $z\in\mathbb{R}^{k}$, we define $\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}$ as the level-set of the map $\xi$. Assume that $\Sigma_{z}$ is non-empty and that the Jacobian of $\xi$ at $x$, denoted by $\nabla\xi(x)\in\mathbb{R}^{d\times k}$, has full rank $k$ at each $x\in\Sigma_{z}$. Under this assumption, the regular value theorem implies that $\Sigma_{z}$ is a (dk)(d-k)-dimensional submanifold of $\mathbb{R}^{d}$; see [16, Chapter 1, Theorem 3.2]. Let $\widetilde{\mu}$ denote the pushforward measure of $\mu$ by the map $\xi$; that is, $\widetilde{\mu}$ is the probability measure of the random variable $\xi(X)\in\mathbb{R}^{k}$ when $X\sim\mu$. Let $\mu_{z}$ denote the conditional probability measure on $\Sigma_{z}$. The law of total expectation implies
$\mathbb{E}_{\mu}(f(X))=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu}\big(f(X)\mid\xi(X)=z\big)\Big]=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu_{z}}\big(f(X)\big)\Big]\,$, (2)
for a test function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, the co-area formula implies that (see [10, Theorem 3.11] and [25, Lemma 3.2])
$\mu_{z}(dx)=\frac{1}{Q(z)}\rho(x)\det\big(\nabla\xi(x)^{\top}\nabla\xi(x)\big)^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(dx)\,,\quad x\in\Sigma_{z}\,$, (3)
and
$\widetilde{\mu}(dz)=Q(z)\,dz\,,\quad z\in\mathbb{R}^{k}\,$, (4)
where $\sigma_{\Sigma_{z}}$ denotes the surface measure of $\Sigma_{z}$ and
$Q(z)=\int_{\Sigma_{z}}\rho(x)\det\big(\nabla\xi(x)^{\top}\nabla\xi(x)\big)^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(dx)\,,\quad z\in\mathbb{R}^{k}\,$, (5)
is a normalizing factor in (3). For processes whose invariant distribution is the Boltzmann–Gibbs distribution $\mu(dx)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}dx$, the free energy associated with the map $\xi$ is defined as
$F(z)=-\beta^{-1}\ln Q(z)\,,\quad z\in\mathbb{R}^{k}\,$, (6)
and, from (4), we have
$\widetilde{\mu}(dz)=\mathrm{e}^{-\beta F(z)}\,dz\,,\quad z\in\mathbb{R}^{k}\,.$ (7)
3 Learning approach
In this section, we present our methodology for learning generative models on the level-sets of a smooth map $\xi$. We first describe the standard approach based on flow-matching for learning the conditional distributions $\mu_{z}$ corresponding to different values $z\in\mathbb{R}^{k}$. Then, we discuss an enhanced learning strategy that leverages biased trajectory data to improve the quality of the learning in regions of low probability under the pushforward distribution $\widetilde{\mu}$. Finally, we explain how such biased trajectory data can be generated in practice using adaptive biasing techniques.
3.1 Standard learning approach
Our goal is to sample the conditional probability measure $\mu_{z}$ on the level-sets of $\xi$ for different values $z\in\mathbb{R}^{k}$. For this purpose, we adopt the flow-matching generative modeling framework [32, 31, 1], which employs ordinary differential equations (ODEs) in $\mathbb{R}^{d}$
$\frac{dX_{t}}{dt}=v(X_{t},t)\,,\quad X_{0}\sim\nu_{\mathrm{prior}}\,,\quad t\in[0,1]$ (8)
to transform a prior distribution $\nu_{\mathrm{prior}}$ to a target distribution $\nu$ that is represented by data. For this purpose, the vector field $v$ in (8) is learned by minimizing the standard flow-matching objective [32]
$\mathcal{L}_{\mathrm{FM},\nu}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\nu}\left\|v\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,.$ (9)
It is known that the minimizer of (9) leads to an ODE (8) that transforms the prior distribution $\nu_{\mathrm{prior}}$ to the target $\nu$.
Adapting this framework to our setting, we propose to learn the ODEs
$\frac{dX_{t}}{dt}=v^{(z)}(X_{t},t)\,,\quad X_{0}\sim\nu_{\mathrm{prior}}\,,\quad t\in[0,1],$ (10)
parametrized by a variable $z\in\mathbb{R}^{k}$, such that, for a fixed $z\in\mathbb{R}^{k}$, the system (10) transforms the prior distribution from $t=0$ to the conditional measure $\mu_{z}$ at $t=1$. To achieve this goal, we consider the following objective for learning the vector field in (10):
$\mathcal{L}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\mu}\left\|v^{(\xi(X_{1}))}\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,$, (11)
where $\mu$ is the probability distribution introduced in the previous section. To justify (11), notice that we can derive
$\mathcal{L}(v)$ $=\mathbb{E}_{X_{1}\sim\mu}\mathbb{E}_{t\sim U([0,1]),X_{0}\sim\nu_{\mathrm{prior}}}\left\|v^{(\xi(X_{1}))}\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}$ (12)
$=\mathbb{E}_{z\sim\widetilde{\mu}}\left[\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\mu_{z}}\left\|v^{(z)}\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\right]$
$=\mathbb{E}_{z\sim\widetilde{\mu}}\left(\mathcal{L}_{\mathrm{FM},\mu_{z}}(v^{(z)})\right)\,$,
where the second equality follows from the law of total expectation in (2). This implies that minimizing the objective (11) is equivalent to minimizing the standard flow-matching objective (9) simultaneously for different target distributions $\mu_{z}$. Therefore, it allows us to learn the ODE (10) for generative modeling on level-sets $\Sigma_{z}$ corresponding to different values $z$ at the same time.
In practice, (11) is replaced by the empirical objective
$\mathcal{L}^{(N)}(v)=\frac{1}{N}\sum_{n=1}^{N}\left\|v^{(\xi(X^{(n)}))}\big((1-t_{n})X_{0}^{(n)}+t_{n}X^{(n)},t\big)-(X^{(n)}-X^{(n)}_{0})\right\|^{2}\,$, (13)
where $t_{1},t_{2},\dots,t_{N}$ are drawn independently from $U([0,1])$, $X_{0}^{(1)},X_{0}^{(1)},\dots,X_{0}^{(N)}$ are drawn independently from $\nu_{\mathrm{prior}}$, and $X^{(1)},X^{(2)},\dots,X^{(N)}$ are the data points.
3.2 Enhanced learning approach with biased trajectory data
We consider the case where the target distribution $\mu$ is the Boltzmann–Gibbs distribution in (1), which is the invariant distribution of a diffusion process $x_{t}$ in $\mathbb{R}^{d}$. For simplicity, we assume that $x_{t}$ is governed by the overdamped Langevin dynamics
$dx_{t}=-\nabla V(x_{t})\,dt+\sqrt{2\beta^{-1}}\,dw_{t}\,,\quad t\geq 0\,$, (14)
where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that increases to infinity sufficiently fast as $|x|\rightarrow+\infty$ and $w_{t}$ is a standard Brownian motion in $\mathbb{R}^{d}$. In this case, it is known that the Boltzmann–Gibbs distribution $\mu$ in (1) is invariant under the process $x_{t}$ in (14) and that $x_{t}$ is ergodic with respect to $\mu$. Thanks to the ergodicity of the process, the empirical objective can be evaluated from trajectory data $X^{(1)},X^{(2)},\dots,X^{(N)}$ obtained by sampling (14) using a numerical scheme at time $t=h,2h,\dots,Nh$, where $h$ denotes the step-size and $N$ is the total number of sampling steps.
We are particularly interested in the case where the potential $V$ has multiple local minima and $\beta$ is (relatively) large. In this case, the probability density of the invariant distribution $\mu$ is highly non-uniform in space, since the probability that $x_{t}$ is within the vicinity of one of the local minima (metastable regions) is much larger than the probability that $x_{t}$ is elsewhere. Likewise, for a proper CV map $\xi$, the pushforward distribution $\widetilde{\mu}$ of $\mu$ defined in (4) is highly non-uniform in $\mathbb{R}^{k}$ as well. In this case, the generative model that is learned by minimizing the objective (11) (or its empirical counterpart (13)) using training data sampled from (14) may be inferior for values $z$ that belong to low-density regions under $\widetilde{\mu}$, due to the relatively small contribution of the loss in such regions to the aggregated objective (11) (see (12)).
In order to overcome the aforementioned issue, we utilize the observation (as can be seen from the derivation (12)) that different probability distributions of $X_{1}$ can be chosen in (11), as long as the corresponding conditional probability distributions coincide with the conditional probability distributions $\mu_{z}$ of $\mu$. In particular, let us consider the reweighted (RW) probability distribution defined by
$\mu_{\mathrm{RW}}(dx)=\bar{Z}^{-1}Z^{-1}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx\,,\quad x\in\mathbb{R}^{d}\,$, (15)
where $A:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\bar{Z}=Z^{-1}\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx$ is a normalizing factor. Then, for each $z\in\mathbb{R}^{k}$, it is straightforward to verify that the conditional probability distribution of $\mu_{\mathrm{RW}}$ on the level-set $\Sigma_{z}$ is $\mu_{z}$ given in (3). Moreover, using (5) and (6), we can obtain that the pushforward of $\mu_{\mathrm{RW}}$ by the map $\xi$ satisfies
$\widetilde{\mu}_{\mathrm{RW}}(dz)=\bar{Z}^{-1}\mathrm{e}^{-\beta\big(F(z)-A(z)\big)}dz\,,\quad z\in\mathbb{R}^{k}\,$, (16)
where $F$ is the free energy (6) associated to the map $\xi$. With the choice of the probability distribution $\mu_{\mathrm{RW}}$ in (15), the objective (11) becomes
$\mathcal{L}_{\mathrm{RW}}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\mu_{\mathrm{RW}}}\left\|v^{(\xi(X_{1}))}\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,$, (17)
and by analogy with (18) we can derive that
$\mathcal{L}_{\mathrm{RW}}(v)=$ $\mathbb{E}_{z\sim\widetilde{\mu}_{\mathrm{RW}}}\Big(\mathcal{L}_{\mathrm{FM},\mu_{z}}(v^{(z)})\Big)\,.$ (18)
From (16), we can see that the density of $\widetilde{\mu}_{\mathrm{RW}}$ is more uniform than the density of $\widetilde{\mu}$ in (7) when $A$ is close to the free energy $F$. Therefore, (18) implies that in this case the contribution of the loss for values $z$ in low-density regions under $\widetilde{\mu}$ is enhanced in the objective (17).
To generate training data, we notice that $\mu_{\mathrm{RW}}$ in (15) is the invariant distribution of the process
$dx_{t}=-\nabla(V-A\circ\xi)(x_{t})\,dt+\sqrt{2\beta^{-1}}\,dw_{t}\,,\quad t\geq 0\,.$ (19)
To ensure that $A$ approximates the free energy $F$, we adopt the adaptive biasing force (ABF) method [7, 14, 11, 6, 30]. Concretely, we generate training data by simulating the biased process
$dx_{t}=-\nabla V(x_{t})\,dt+\nabla\xi(x_{t})f_{t}(\xi(x_{t}))\,dt+\sqrt{2\beta^{-1}}\,dw_{t}\,,\quad t\geq 0\,$, (20)
where $f_{t}(z)$ is an estimation of the gradient of the free energy $F$ (i.e. the mean force), based on the expression [25, Lemma 3.9]:
$\frac{\partial F}{\partial z_{i}}(z)=\mathbb{E}_{\mu_{z}}\left[\sum_{l=1}^{d}\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\frac{\partial V}{\partial x_{l}}-\beta^{-1}\sum_{l=1}^{d}\frac{\partial}{\partial x_{l}}\bigg(\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\bigg)\right]\,,\quad i=1,2,\dots,k\,$, (21)
where $(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}$ denotes the $ij$-component of the inverse of the matrix $\nabla\xi^{\top}\nabla\xi\in\mathbb{R}^{k\times k}$. In practice, a grid in $\mathbb{R}^{k}$ is defined and the mean value (21) is estimated for each cell of the grid. In each simulation step, the estimation is updated for the cell which the current state $\xi(x_{t})$ falls into. When the cell has been visited for more than a certain (pre-defined) number of times, the current value of the estimation is used as the biasing term $f_{t}(\xi(x_{t}))$ in (20). As $t\rightarrow+\infty$, we can expect that the empirical estimations converge to the true means in (21) and, accordingly, the process (20) converges to its equilibrium counterpart (19) with $A=F$.
3.3 Projection step
Once the model is learned using the approach described in the previous subsection, it can be used to generate new data points on different target level-sets of $\xi$ by integrating the ODE (10) with different values $z\in\mathbb{R}^{k}$. However, due to numerical errors, the generated data may not lie exactly on the target level-set. Here, we propose to implement an additional post-processing step optionally in order to project the generated data onto the target level-set. Also refer to [33] for discussions on projection onto manifolds in the context of generative modeling. The first method is the projection to the closest point on the level-set, which amounts to solving
$\bar{x}:=\arg\min_{y\in\Sigma_{z}}\|x-y\|^{2}$ (22)
for each generated point $x$. However, unless the level-set $\Sigma_{z}$ can be parametrized by certain coordinates, solving (22) is computationally infeasible because it is a constrained (high-dimensional) optimization task. Another simpler method is to project the generated state $x$ along the ODE flow [49]
$\frac{dY_{t}}{dt}=-\nabla G(Y_{t})\,,\quad Y_{0}=x\,$, (23)
where the function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined as $G(y):=\frac{1}{2}\|\xi(y)-z\|^{2}$ for $y\in\mathbb{R}^{d}$. Thanks to the gradient structure, it can be shown that, under certain regularity assumptions on $\xi$, the quantity $G(Y_{t})$ decreases to zero exponentially fast as $t\rightarrow+\infty$, implying that the state evolved by the flow of (23) approaches the level-set $\Sigma_{z}$. In practice, to perform the projection, it suffices to integrate the ODE (23) up to a time when certain criteria is met, i.e. $\|\xi(Y_{t})-z\|<\epsilon_{\mathrm{tol}}$, where $\epsilon_{\mathrm{tol}}>0$ is a small parameter specifying the tolerance.
4 Experiments
In this section, we demonstrate the capability of our approach through different representative datasets. The section is organized as follows: we first consider a synthetic dataset in $\mathbb{R}^{2}$ to illustrate the method. Next, we evaluate its performance on the trajectory data of the system under the Müller–Brown potential, a widely used benchmark system exhibiting metastability. Finally, we apply the approach to the alanine dipeptide molecular system, showcasing its applicability to molecular systems with complex energy landscapes. The neural network and the training were implemented using PyTorch [39].
4.1 2D Synthetic Dataset
Figure 1: (a) Scatter plot of the original 2D dataset sampled using the make_circles function from scikit-learn [40]. (b) Density estimate of the corresponding CV map $\xi(x)=x_{1}^{2}+x_{2}^{2}$.
Figure 2: Results for the 2D dataset. (a) Mean value of $\xi$ of generated samples compared to the intended CV value. (b) Mean deviation of $\xi$ on generated samples from the target CV value. (c) Proportion of samples with positive $x_{1}$ values in the original and generated samples for different target CV values.
Figure 3: Evolution of the two-dimensional samples generated by the conditional ODE flow for a fixed target CV value $z=0.6$. Each panel shows the particle positions at different integration times $t$, illustrating how the initial Gaussian cloud progressively morphs into the target distribution.
The 2D dataset (Figure 1(a)) is sampled using the make_circles function from the package scikit-learn [40]. The CV map is defined as $\xi(x)=x_{1}^{2}+x_{2}^{2}$ for $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, whose empirical density is shown in Figure 1(b).
The prior $\nu_{\mathrm{prior}}$ is chosen as the standard Gaussian distribution in $\mathbb{R}^{2}$. The vector field $v$ in (10) is parameterized as a fully connected feedforward neural network, comprising a total of four layers: an input layer, two hidden layers with $128$ neurons each and Tanh activations, followed by an output layer of size $2$. The input consists of the spatial coordinate $x$ in $\mathbb{R}^{2}$, a time scalar $t$, and the target CV value $z$ (with dimension 1). To learn the vector field $v$, the empirical objective $\mathcal{L}^{(N)}(v)$ in (13) is minimized using the Adam optimizer for up to 1000 epochs with early stopping (patience of 200), a learning rate of $10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $1000$. For other parameters, their default values in PyTorch were used.
Next, we evaluate the performance of the learned model for generating states conditioned on the value $z$. To this end, we choose the values of $z$ uniformly within an extended range around the values visited in the dataset (with a 5% margin), and for each chosen value of $z$, $1000$ new states are generated by integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $1000$ time steps, starting from the standard Gaussian prior at $t=0$. Based on the generated states, we estimate the mean value of $\xi$, the mean deviation of $\xi$ from the target CV value $z$, defined as
$\mathrm{Deviation}(z)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\Big(\xi(X^{(i)})-z\Big)^{2}},$ (24)
where $(X^{(i)})_{1\leq i\leq N}$ are the generated samples, as well as the proportion of states with $x_{1}\geq 0$, which are shown in Figures 2(a)2(b), and 2(c), respectively. The results in Figure 2 demonstrate that the proposed model performs consistently well across all evaluation metrics in the two-dimensional case. The generated CV values closely match the intended targets, the deviation remains low across the entire range of values, and the model accurately learns the conditional distribution.
Finally, Figure 3 presents the temporal evolution of the generated distribution for $z=0.6$, obtained by integrating the ODE from a standard Gaussian initial cloud of $2000$ points. Successive snapshots are taken at increasing integration times $t\in\{0,0.25,0.50,0.75,1.0\}$, showing how the initially isotropic Gaussian distribution gradually transforms into the target ring-shaped configuration.
4.2 Müller–Brown Dataset
We apply our approach to the two-dimensional Boltzmann–Gibbs distribution (1) with the Müller–Brown potential [38]. As shown in Figure 4, the potential has two deep wells (i.e. low-potential regions), separated by a shallow well. At low temperatures, the corresponding dynamics exhibits metastability with a curved transition pathway, making it an ideal model system for studying the complex behavior of metastable dynamics.
Model Description and Training.
The dataset is prepared by sampling the trajectory of the overdamped Langevin dynamics (14) up to time $t=600$ (corresponding to $3\times 10^{6}$ steps), with the integration time step $\Delta t=2\times 10^{-4}$ and inverse temperature $\beta=0.1$. The starting point is chosen in the deepest well, with configuration (0.6,1.2)(-0.6,1.2). By recording states every 100 steps, we obtained a dataset consisting of $3\times 10^{4}$ configurations.
Figure 4: Heatmap and contour representation of the Müller–Brown potential landscape. The two major low-potential regions and the shallow low-potential region are shown in dark blue and light blue, respectively, while the contour lines highlight specific energy levels.
In order to define a CV map, we first trained an autoencoder with a one-dimensional bottleneck, modeled as a fully connected feedforward neural network, by minimizing the mean squared reconstruction error on the generated trajectory data. The encoder network consists of an input layer of size 2, three hidden layers of size 32 with Tanh activation, and an output layer of size 1. The decoder network consists of an input layer of size 1, a single hidden layer of size 32 with Tanh activation, and an output layer of size 2. Training was performed for $200$ epochs using the Adam optimizer, with a learning rate of $10^{-3}$ and a batch size of $128$. The CV map $\xi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined by the trained encoder. Figure 5(a) shows the level-sets of $\xi$ with the generated trajectory data superimposed, whereas Figure 5(b) shows the scatter plot of the potential energy $V(x)$ for each state $x$ in the trajectory against the encoded CV value $\xi(x)$.
To test the enhanced learning technique discussed in Section 3.2, we also employed the ABF dynamics (20) with the learned CV map $\xi$, where the biasing force $f_{t}$ was computed based on (21) on a uniform grid with grid spacing $0.1$ that covers the range [2.6,3.5][-2.6,3.5]. To ensure accurate estimation, the biasing force in a cell of the grid is applied to the system only after 100 samples have been used to estimate (21). The biased simulation was run for $1.5\times 10^{5}$ steps and, after the initial $5\times 10^{4}$ simulation steps (equilibration), states were recorded every $10$ steps, resulting in a trajectory dataset of size $10^{4}$. Similarly to Figures 5(a)5(b), Figure 6(a) and 6(b) show the scatter plot of the biased trajectory data superimposed on the level-sets of $\xi$ and the scatter plot of potential values $V(x)$ against the encoded CV value $\xi(x)$, respectively.
We trained our generative model (10) using the generated trajectory data and the learned CV map $\xi$. The vector field $v$ is modeled by a fully connected feedforward neural network, which consists of an input layer of size 4 (i.e. the total dimension of the state $x\in\mathbb{R}^{2}$, time $t\in[0,1]$ and the target CV value $z\in\mathbb{R}$), three hidden layers of size 128 with Tanh activation, and an output layer of size $2$. Training was performed for 3000 epochs by minimizing the objective (13) (respectively, the empirical counterpart of (17) for biased trajectory data) using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $512$.
Results.
We evaluated the trained generative models by generating new states with prescribed target CV values 2.0,0.0-2.0,0.0 and $2.0$. For each target CV value $z$, $10^{3}$ states sampled from the two-dimensional standard Gaussian distribution were propagated by numerically integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $100$ uniform time steps. The resulting final states at time $t=1$ were interpreted as samples associated with the chosen CV value $z$. These samples are visualized in Figures 7(a) and 8(a). Comparing these two figures reveals that the model trained using biased trajectory data produces spatially consistent samples distributed along the correct CV contours (Figure 8(a)), whereas the model trained using unbiased trajectory data tends to generate inconsistent samples when the target level-sets are not well covered by the training data (i.e. level-sets corresponding to z=2.0z=-2.0 and $z=0.0$ in Figure 7(a)). To quantify the evaluation, the encoded CVs of generated samples were compared to the target value $z$, and the root mean squared deviation defined in (24) was computed for a uniform grid of target values $z\in[-2.9,3.6]$. The deviation curves in Figure 9 summarize this evaluation: the model trained on biased trajectories achieves a lower deviation across the CV range. Figure 10 shows snapshots of the temporal evolution of the point distributions under the learned flow model (10), taken at specific time points $t=0.0,0.5,0.9,0.95,1.0$, illustrating the transformation from the initial standard Gaussian distribution to the final distribution conditioned on the latent variable $z$.
Figure 5: (a) Trajectory samples overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value.
Figure 6: (a) Biased trajectory samples obtained from ABF simulation overlaid on level-sets of the learned CV map $\xi$. The color bar indicates the corresponding CV values across the configuration space. (b) Scatter plot of the potential energy versus the CV value for states in the biased trajectory.
Figure 7: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on unbiased trajectory data. (a) Without projection. (b) With projection.
Figure 8: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on biased trajectory data. (a) Without projection. (b) With projection.
Figure 9: Comparison of the deviations defined in (24) for different target CV values between the model trained on unbiased trajectory data and the model trained on biased trajectory data (ABF).
We also tested the method introduced in Section 3.3 for projecting the generated samples. For each target CV value $z$, the samples obtained from the generative models (as shown in Figures 7(a) and 8(a)) were propagated by numerically integrating the ODE 23 using the explicit Euler scheme, where we used step size $0.01$, the maximal number of integration steps $7000$ and tolerance $\epsilon_{\mathrm{tol}}=10^{-3}$. The resulting states are shown in Figures 7(b) and 8(b), respectively, from which we can conclude that the generated samples were successfully projected onto the corresponding level-sets.
Finally, we assessed the closeness between the empirical distribution of the generated samples and the conditional distribution on the corresponding level-set. We chose z=2.0z=-2.0 and sampled $5\times 10^{4}$ states on the corresponding level-set using the constrained sampling scheme in [5] with a step size $2.0\times 10^{-4}$. Each state $x$ was assigned the weight $|\nabla\xi(x)|^{-1}$, which is proportional to the likelihood ratio between the conditional distribution (3) and the invariant distribution of the constrained numerical scheme in the continuous-time limit, so that the weighted samples can be used to approximate the true conditional distribution on the level-set (see [5, 49] for detailed discussions). With these data, we estimated the densities of the first component $x_{1}$ for the samples generated by the learned generative model, their projections on the level-set by the method in Section 3.3, and the weighted samples obtained by constrained sampling. The results are shown in Figure 11(a) and Figure 11(b) for the model trained using unbiased trajectory data and the model trained using biased trajectory data, respectively. From these two figures, we can conclude that the projection step maps the generated states onto the level-set without modifying their distribution. Moreover, it can be observed that, compared to the model learned using unbiased trajectory data, the generative model learned using biased trajectory generates data whose distribution is closer to the true conditional distribution. We also computed probability densities of other quantities such as the potential energy and distance to the mean. The results are similar to those in Figure 11 and therefore are not shown here.
Figure 10: Time evolution of the data points under the learned model (10) for three different target values of $z\in\{-2.0,0.0,2.0\}$. The top part of the figure corresponds to the model trained with unbiased trajectory data, and the bottom part corresponds to the model trained with biased trajectory data. Contours and filled colormaps show the level-sets and the profile of the CV map $\xi$, respectively. Points corresponding to each $z$ value are drawn in distinct colors to track their evolution over time.
Figure 11: Density estimations of the first component $x_{1}$ for the samples generated using the generative model for the value z=2.0z=-2.0 (without projection), their projections onto the level-sets (with projection), and the samples obtained by constrained sampling on the level-set (reference). (a) Results for the generative model trained using unbiased trajectory data. (b) Results for the generative model trained using biased trajectory data.
4.3 Alanine Dipeptide
Model Description and Training.
In this study, we applied our generative approach to alanine dipeptide, which is a small benchmark molecular system consisting of 22 atoms. Our goal is to train the ODE model (10) for generative modeling of configurations of the system’s 10 non-Hydrogen atoms conditioned on its two backbone dihedral angles $\phi$ and $\psi$.
Trajectory data of the alanine dipeptide system were generated by performing molecular dynamics simulations for $t=1500\,\textrm{ns}$ using the GROMACS package [47], where we adopted the same simulation parameters as in the previous work [24] and the states were stored every $10$ ps, resulting in a trajectory dataset of $1.5\times 10^{5}$ states. Only Cartesian coordinates of the 10 non-Hydrogen atoms were then extracted using MDAnalysis [12, 37] and aligned to a reference configuration via the Kabsch algorithm [20], thereby eliminating global translational and rotational motions. The aligned coordinates were then standardized by subtracting the empirical mean and dividing by the empirical standard deviation before being used as training data. Figure 12(a) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles.
In addition to the unbiased MD data, we also generated an independent trajectory by performing an ABF simulation for $t=500\,\textrm{ns}$ along the backbone dihedral angles $\phi$ and $\psi$. The ABF method applies a time-dependent biasing force to enhance the sampling of rarely visited regions of conformational space. The simulation was performed using the ColVars module [11] and the GROMACS package. The states were stored every $10$ steps and the resulting trajectory dataset consists of $5\times 10^{4}$ states. Figure 12(b) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles. All ABF-biased coordinates were preprocessed in the same manner as the unbiased data—centered, aligned to a reference structure, flattened, and then standardized.
For the modeling of the vector field $v$ in (10), we implemented a fully connected feedforward neural network that takes as input the atomic coordinates $x\in\mathbb{R}^{30}$, the temporal parameter $t\in[0,1]$, and the CV value $z\in\mathbb{R}^{2}$, which correspond to the standardized values of $(\phi,\psi)$. The feedforward network consists of five hidden layers with 512 hidden units and Tanh activations, while the output layer returns the vector field in $\mathbb{R}^{30}$.
The training of the vector field $v$ was performed by minimizing the objective (13) for 1000 epochs, using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of 512, and default PyTorch parameters (e.g., for momentum coefficients, numerical stability parameter, and weight decay). After training, new samples in $\mathbb{R}^{30}$ can be generated conditioned on the value of $z$ (i.e. the target value of $(\phi,\psi)$) by integrating the ODE (10) from random initial positions sampled from the standard Gaussian distribution. The generated samples are then denormalized and reshaped to obtain three-dimensional coordinates of the $10$ non-Hydrogen atoms.
Results.
To assess the quality of the generated conformations, we chose the target dihedral angles $(\phi,\psi)\in\{(-75^{\circ},150^{\circ}),(52.49^{\circ},34.95^{\circ}),(57.5^{\circ},-126.78^{\circ})\}$, which correspond respectively to the densely populated region, a secondary metastable region, and an isolated point in a low-population region in the dihedral angle space (see Figure 12(a)). For each of these values $(\phi,\psi)$, we generated 1000 samples and identified the (reference) configuration in the trajectory whose dihedral angles are closest to $(\phi,\psi)$. Root mean square deviation (RMSD) values of the generated samples from the reference were computed after Kabsch alignment. Histograms of RMSD values were constructed with fixed bin widths ($0.02$ Å) to visualize the distribution of deviations in different regions represented by these values of dihedral angles. These distributions are shown for the model trained using the unbiased dataset in Figure 13 and for the model trained using the biased dataset in Figure 14, respectively.
RMSD distributions between nearby real structures are shown for the unbiased trajectory dataset in Figures 15(a) and 15(b), serving as a baseline for intrinsic structural variability. In the left panel (Figure 15(a)), all pairs of configurations whose dihedral angles $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$ satisfy $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1^{\circ}$ were identified using a nearest neighbors search in the dihedral angle space. For each pair, the RMSD was computed after Kabsch alignment to remove global translations and rotations. The right panel (Figure 15(b)) extends this analysis by binning the $\phi$ and $\psi$ angles of configurations into a $50\times 50$ grid and computing the mean RMSD of all conformations within each bin (empty bins were ignored). For the ABF-biased dataset, the same evaluation workflow was applied, and the results were nearly identical to those for the unbiased dataset. Comparing the histograms in Figure 13 and Figure 15, we can conclude that the mean deviations predicted by the trained model are of the same order as the intrinsic structural variability observed in the unbiased dataset.
To assess the angular consistency of the generated configurations across the dihedral space, we computed heatmaps of mean deviations (see (24)) per $(\phi,\psi)$-bin for the generative models trained using both the unbiased and the ABF-biased datasets, as shown in Figure 16. For this analysis, the $\phi$ and $\psi$ angles in the range $180^{\circ},180^{\circ}]$ were binned into intervals of equal length $20^{\circ}$, and the centers of each bin were used as target CV values. For each target CV value, $N=200$ configurations were generated using the trained ODE models. The resulting mean deviations are shown in Figure 16(a) for the model trained using the unbiased trajectory data, where small deviations are concentrated in the high-probability regions of the CV space, and in Figure 16(b) for the model trained using the ABF-biased trajectory data, which exhibit uniformly low deviations across the full CV space.
Similarly to the previous example, the projection method introduced in Section 3.3 can be applied as a post-processing step to modify the generated configurations so that they satisfy the constraints on the target dihedral angles. We refer to the previous work [34, Appendices B.2 and B.4] for related numerical study.
Figure 12: Empirical densities of the dihedral angles of the trajectory data. (a) Trajectory from unbiased simulation. (b) Trajectory from ABF simulation.
Figure 13: Results of the generative model trained with the unbiased dataset. Distributions of RMSD values (in Å) for generated samples at three representative CV locations: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$ in a well-populated region (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$ in the secondary metastable basin, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$ corresponding to an isolated point in the low-population region.
Figure 14: Results of the generative model trained with the ABF-biased dataset. Distributions of RMSD values (in Å) for generated samples at the same representative CV locations as in Figure 13: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$, (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$.
Figure 15: Analysis of intrinsic structural variability in the unbiased real dataset. (a) Distribution of RMSD values (in Å) between pairs of nearby configurations satisfying $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1$. (b) Histogram of mean RMSD values computed for each $50\times 50$ $(\phi,\psi)$-bin in the dihedral angle space, capturing spatial variations in local conformational variability.
Figure 16: Heat maps of the mean deviation between the target CV value $(\phi,\psi)$ and the dihedral angles of generated configurations. (a) Unbiased molecular dynamics data. (b) ABF-biased data. Each pixel represents the mean circular distance in degrees, quantifying the accuracy with which the conditional generative model reproduces the desired angular constraints. Lower deviation indicates higher fidelity of the generated ensemble to the target CVs.
5 Conclusion
In this work, we have introduced a general framework for the generative modeling of conditional distributions on level-sets of CVs using flow-matching neural ODEs. By conditioning the learned flows on CV values, the model can generate samples across multiple level-sets without requiring explicit geometric information. To further enhance accuracy on level-sets in low-probability regions, we proposed to leverage biased trajectory data from enhanced sampling techniques such as ABF. Validation on concrete datasets demonstrates that our approach accurately reproduces conditional distributions and successfully exploits biased data to improve learning in scarcely sampled regions.
Beyond these results, this study opens several promising directions for future work. First, to better assess the effectiveness of the approach, it is meaningful to study the accuracy of the generative model by comparing certain distances between the generated distributions and the conditional distributions on level-sets. Second, a natural extension would be to combine the approach proposed in this work with methods that propagate the effective (reduced) dynamics in the CV space [21, 50]. Such a combination would lead to generative modeling approaches that can generate new trajectories of stochastic processes while taking into account the essential dynamics.
Overall, our study provides a general and practical approach for conditional generative modeling on manifolds defined by CVs, offering both methodological flexibility and the potential for applications in molecular modeling and beyond.
6 Acknowledgements
Fatima-Zahrae Akhyar acknowledges the Zuse institute Berlin for hosting her internship during which this research was conducted. Wei Zhang has been funded through DFG’s “Eigene Stelle” program (project 524086759) and has been supported by the DFG-grant CRC 1114 “Scaling Cascades in Complex Systems” Project Number 235221301. Gabriel Stoltz was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project EMC2, grant agreement No 810367), and was supported by Hi! PARIS and ANR/France 2030 program (ANR-23-IACL-0005). Christof Schütte has been supported through DFG-grant CRC 1114 (Project A04 “Efficient calculation of slow and stationary scales in molecular dynamics”, Project A05 “Probing scales in equilibrated systems by optimal nonequilibrium forcing”, Project B03 “Multilevel coarse graining of multiscale problems”, and Project C03 “Multiscale modelling and simulation for spatiotemporal master equation”).

### v3

@@
 Abstract
-Given a probability distribution $\mu$ in $\mathbb{R}^{d}$ represented by data, we study in this paper the generative modeling of its conditional probability distributions on the level-sets of a collective variable $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$.
-We propose a general and efficient learning approach that is able to learn generative models on different level-sets of $\xi$ simultaneously.
-To improve the learning quality on level-sets in low-probability regions, we also propose a strategy for data enrichment by utilizing data from enhanced sampling techniques.
+Given a probability distribution $\mu$ in $\mathbb{R}^{d}$ represented by data, we study in this paper the generative modeling of the corresponding conditional probability distributions on the level-sets of a collective variable $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$, where $1\leq k<d$.
+We propose a general and efficient learning approach that can learn generative models on different level-sets of $\xi$ simultaneously.
+To improve the learning quality on level-sets in low-probability regions, we also propose a data enrichment strategy by utilizing data from enhanced sampling techniques.
 We demonstrate the effectiveness of our proposed learning approach through concrete numerical examples.
-The proposed approach is potentially useful for the generative modeling of molecular systems in biophysics, for instance.
+The proposed approach is potentially useful for the generative modeling of molecular systems in biophysics.
 Keywords— collective variables, level-sets, conditional probability distributions, generative modeling, flow-matching
@@
 High-dimensional stochastic processes are ubiquitous in broad research areas such as biophysics, materials science, and climate modeling.
-Understanding the complex dynamical behavior of such processes often requires projecting the system down to a (much) lower-dimensional space via a few features, i.e.
+Understanding the complex dynamical behavior of such processes often requires projecting the system down to a (much) lower-dimensional space defined by a few features, i.e.
 collective variables (CVs) or reaction coordinates, of the system.
-In molecular dynamics, for instance, various enhanced sampling techniques have been developed to overcome the sampling difficulty due to the system’s metastability and to construct the free energy surface in the space of CVs [15].
+In molecular dynamics, for instance, various enhanced sampling techniques have been developed to increase the sampling efficiency and to construct the free energy surface in the space of CVs [15].
 A number of analytical and computational approaches have also been proposed for constructing simpler, lower-dimensional surrogate models associated to a chosen set of CVs of the original high-dimensional system.
-One notable approach is the effective dynamics constructed using conditional expectations, where the coefficients (i.e.
+A notable approach is the effective dynamics constructed using conditional expectations, where the coefficients (i.e.
 drift and diffusion) in the surrogate model are computed by taking averages with respect to the so-called conditional probability distribution on the corresponding level-set of the CVs [21].
-This approach has attracted considerable research interest in the mathematical community [50, 51, 22, 29, 23, 9, 52], contributing to a deeper understanding of the quality of the effective dynamics in approximating the original high-dimensional processes.
-A related research topic in molecular dynamics is sampling on the level-sets of CVs [5, 13, 26], which is useful in both of the aforementioned tasks, i.e.
-free energy computation and computing the coefficients in the effective dynamics.
+This approach has attracted considerable research interest in the mathematical community [50, 51, 22, 29, 23, 9, 52], contributing to a deep understanding on the quality of the effective dynamics in approximating the original high-dimensional processes.
+A common task in both of the aforementioned research directions, i.e.
+free energy calculation and the computation of coefficients in the effective dynamics, is sampling on the level-sets of CVs [5, 13, 26].
 Metropolis-Hastings Monte Carlo methods have been developed for unbiased sampling on submanifolds defined by the level-sets of CVs [48, 27, 28].
-Given a (Boltzmann–Gibbs) distribution in the full space, sampling methods for its conditional distribution on the level-sets of CVs have been developed by constructing ergodic diffusion processes on the level-sets [49, 42].
+Given a (Boltzmann–Gibbs) distribution in the full space, sampling methods for the corresponding conditional distribution on the level-sets of CVs have been developed by constructing ergodic diffusion processes on the level-sets [49, 42].
 More broadly, methods for sampling under constraints have also been studied in computational statistics [46, 35].
@@
 Motivated by these achievements, there has been growing research interest in extending generative modeling approaches for Euclidean data to manifold data [36, 41, 2, 8, 18, 19, 4, 34, 33].
-However, to the best of our knowledge, most of these existing approaches rely on substantial geometric information such as geodesics [8], heat kernel, or eigenfunctions [4], which consequently limits their applicability to manifolds where such information is explicitly known.
+Most of these existing approaches rely on substantial geometric information such as geodesics [8], heat kernel, or eigenfunctions [4].
+As a result, their applicability is limited to manifolds where such information is explicitly known.
 Only the methods proposed in [34, 33] can be applied to manifolds that are implicitly defined as a level-set of a certain function.
-All these approaches are developed for generative modeling on a single and fixed manifold.
-In this paper, we study the generative modeling of conditional probability distributions on the level-sets of CVs and propose a versatile framework for this task.
-Our approach builds upon the flow-matching model [31, 32, 1], in which a neural ordinary differential equation (ODE) is learned to transport samples from a base Gaussian distribution to the target conditional distribution on each level-set of a CV map $\xi$.
-Our method does not require explicit geometric information of the underlying manifolds and therefore is broadly applicable to a wide range of manifolds implicitly defined by CVs.
-Moreover, by varying the target CV value, the method can generate samples across different level-sets of $\xi$, rather than being restricted to a single level-set.
-To further improve learning in regions of low probability density, we introduce a data enrichment strategy that leverages data obtained from enhanced sampling techniques such as adaptive biasing force (ABF) simulations.
+However, to the best of our knowledge, all these approaches are developed for generative modeling on a single and fixed manifold.
+In this paper, we propose an effective generative modeling approach for the conditional probability distributions on the level-sets of CVs.
+Our approach builds upon the flow-matching model [31, 32, 1], in which a neural ordinary differential equation (ODE) is learned to transport a prior (Gaussian) distribution to the target conditional distributions on the level-sets of a CV map $\xi$.
+Our method does not require explicit geometric information of the underlying manifolds and is broadly applicable to a wide range of manifolds that are implicitly defined by CVs.
+By varying the target CV value, the method can generate samples across different level-sets of $\xi$, rather than being restricted to a single level-set.
+To further improve the learning in regions of low probability density, we introduce a data enrichment strategy that leverages data obtained from enhanced sampling techniques such as adaptive biasing force (ABF) simulations.
 The approach is validated on a hierarchy of systems of increasing complexity, from synthetic low-dimensional examples to high-dimensional molecular datasets.
@@
   3. (iii)
-    comprehensive numerical validations that demonstrate the flexibility, robustness, and practical applicability of the method to molecular systems.
+    Comprehensive numerical validations that demonstrate the flexibility, robustness, and practical applicability of the method.
 The remainder of the paper is organized as follows.
@@
 2 Background and notation
-The aim of this section is to introduce the probability distributions of interest in this work, in particular the conditional distribution on the level-sets of a given smooth map, which will play a central role in the following.
-Consider a target distribution $\mu$ over $\mathbb{R}^{d}$ with a positive probability density $\rho$ with respect to the Lebesgue measure, namely $\mu(dx)=\rho(x)\,dx$.
+The aim of this section is to introduce the probability distributions of interest in this work, in particular the conditional distributions on the level-sets of a given smooth map.
+Consider a distribution $\mu$ over $\mathbb{R}^{d}$ with a positive probability density $\rho$ with respect to the Lebesgue measure, namely $\mu(dx)=\rho(x)\,dx$.
 Assume that a $C^{2}$-smooth map $\xi:\mathbb{R}^{d}\rightarrow\mathbb{R}^{k}$ is given, where $1\leq k<d$.
-We are particularly interested in applications in statistical physics, where $\mu$ arises as the invariant distribution of certain ergodic diffusion processes $x_{t}$ in $\mathbb{R}^{d}$ and $\xi$ corresponds to the CV of the system.
+We are particularly interested in applications in statistical physics, where $\mu$ arises as the invariant distribution of certain ergodic diffusion process $x_{t}$ in $\mathbb{R}^{d}$ and $\xi$ corresponds to the CV of the system.
 In this case, we often refer to the Boltzmann–Gibbs distribution, whose density with respect to the Lebesgue measure is
 $\rho(x)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)},\quad x\in\mathbb{R}^{d}\,$, (1)
-where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that grows sufficiently fast to infinity as $|x|\rightarrow+\infty$, $\beta$ is a positive constant whose inverse is proportional to the system’s temperature, and $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,dx$ is the normalizing constant.
-We introduce a few quantities related to the measure $\mu$ and the map $\xi$.
+where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that grows sufficiently fast to infinity as $|x|\rightarrow+\infty$, $\beta$ is a positive constant whose inverse is proportional to the system’s temperature, and $Z=\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta V(x)}\,dx$ is the normalizing constant that we assume to be finite.
+Below we introduce a few quantities related to the measure $\mu$ and the map $\xi$.
 Further discussions can be found in [25, Section 3.2.1] and [50].
-Given $z\in\mathbb{R}^{k}$, we define $\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}$ as the level-set of the map $\xi$.
-Assume that $\Sigma_{z}$ is non-empty and that the Jacobian of $\xi$ at $x$, denoted by $\nabla\xi(x)\in\mathbb{R}^{d\times k}$, has full rank $k$ at each $x\in\Sigma_{z}$.
-Under this assumption, the regular value theorem implies that $\Sigma_{z}$ is a (dk)(d-k)-dimensional submanifold of $\mathbb{R}^{d}$; see [16, Chapter 1, Theorem 3.2].
-Let $\widetilde{\mu}$ denote the pushforward measure of $\mu$ by the map $\xi$; that is, $\widetilde{\mu}$ is the probability measure of the random variable $\xi(X)\in\mathbb{R}^{k}$ when $X\sim\mu$.
-Let $\mu_{z}$ denote the conditional probability measure on $\Sigma_{z}$.
+Given $z\in\mathbb{R}^{k}$, the level-set of the map $\xi$ corresponding to $z$ is defined as
+$\Sigma_{z}:=\{x\in\mathbb{R}^{d}\,|\,\xi(x)=z\}.$
+Assume that $\Sigma_{z}$ is non-empty and that $\nabla\xi(x)\in\mathbb{R}^{d\times k}$, the Jacobian of $\xi$, has full rank $k$ at each $x\in\Sigma_{z}$.
+Under these assumptions, the regular value theorem asserts that $\Sigma_{z}$ is a (dk)(d-k)-dimensional submanifold of $\mathbb{R}^{d}$; see [16, Chapter 1, Theorem 3.2].
+Let $\widetilde{\mu}$ denote the pushforward measure of $\mu$ by the map $\xi$; that is, $\widetilde{\mu}$ is the probability distribution of the random variable $\xi(X)\in\mathbb{R}^{k}$ when $X\sim\mu$.
+Let $\mu_{z}$ denote the conditional probability distribution on $\Sigma_{z}$, i.e.
+the probability distribution of $X$ conditioned on the event that $\xi(X)=z$.
 The law of total expectation implies
-$\mathbb{E}_{\mu}(f(X))=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu}\big(f(X)\mid\xi(X)=z\big)\Big]=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu_{z}}\big(f(X)\big)\Big]\,$, (2)
-for a test function $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, the co-area formula implies that (see [10, Theorem 3.11] and [25, Lemma 3.2])
+$\mathbb{E}_{X\sim\mu}(f(X))=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu}\Big(f(X)\,\Big|\,\xi(X)=z\Big)\Big]=\mathbb{E}_{z\sim\widetilde{\mu}}\Big[\mathbb{E}_{X\sim\mu_{z}}\big(f(X)\big)\Big]\,$, (2)
+for test functions $f:\mathbb{R}^{d}\rightarrow\mathbb{R}$. Moreover, the co-area formula implies that (see [10, Theorem 3.11] and [25, Lemma 3.2])
 $\mu_{z}(dx)=\frac{1}{Q(z)}\rho(x)\det\big(\nabla\xi(x)^{\top}\nabla\xi(x)\big)^{-\frac{1}{2}}\sigma_{\Sigma_{z}}(dx)\,,\quad x\in\Sigma_{z}\,$, (3)
@@
 is a normalizing factor in (3).
-For processes whose invariant distribution is the Boltzmann–Gibbs distribution $\mu(dx)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}dx$, the free energy associated with the map $\xi$ is defined as
+For processes whose invariant distribution is the Boltzmann–Gibbs distribution $\mu(dx)=\frac{1}{Z}\mathrm{e}^{-\beta V(x)}dx$, the free energy associated to the map $\xi$ is defined as
 $F(z)=-\beta^{-1}\ln Q(z)\,,\quad z\in\mathbb{R}^{k}\,$, (6)
@@
 We first describe the standard approach based on flow-matching for learning the conditional distributions $\mu_{z}$ corresponding to different values $z\in\mathbb{R}^{k}$.
-Then, we discuss an enhanced learning strategy that leverages biased trajectory data to improve the quality of the learning in regions of low probability under the pushforward distribution $\widetilde{\mu}$.
+Then, we propose an enhanced learning strategy that leverages biased trajectory data to improve the quality of the learning in regions of low probability under the pushforward distribution $\widetilde{\mu}$.
 Finally, we explain how such biased trajectory data can be generated in practice using adaptive biasing techniques.
 3.1 Standard learning approach
-Our goal is to sample the conditional probability measure $\mu_{z}$ on the level-sets of $\xi$ for different values $z\in\mathbb{R}^{k}$.
-For this purpose, we adopt the flow-matching generative modeling framework [32, 31, 1], which employs ordinary differential equations (ODEs) in $\mathbb{R}^{d}$
+Our goal is to sample the conditional probability distributions $\mu_{z}$ for different values $z\in\mathbb{R}^{k}$.
+For this purpose, we adopt the flow-matching generative modeling framework [32, 31, 1], which employs the ordinary differential equation (ODE) in $\mathbb{R}^{d}$
 $\frac{dX_{t}}{dt}=v(X_{t},t)\,,\quad X_{0}\sim\nu_{\mathrm{prior}}\,,\quad t\in[0,1]$ (8)
@@
 For this purpose, the vector field $v$ in (8) is learned by minimizing the standard flow-matching objective [32]
-$\mathcal{L}_{\mathrm{FM},\nu}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\nu}\left\|v\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,.$ (9)
+$\mathcal{L}_{\mathrm{FM},\nu}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\nu}\left\|v\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,$, (9)
+where $U([0,1])$ denotes the uniform distribution on $[0,1]$.
 It is known that the minimizer of (9) leads to an ODE (8) that transforms the prior distribution $\nu_{\mathrm{prior}}$ to the target $\nu$.
-Adapting this framework to our setting, we propose to learn the ODEs
+Adapting this framework to our setting, we propose to learn the ODE
 $\frac{dX_{t}}{dt}=v^{(z)}(X_{t},t)\,,\quad X_{0}\sim\nu_{\mathrm{prior}}\,,\quad t\in[0,1],$ (10)
-parametrized by a variable $z\in\mathbb{R}^{k}$, such that, for a fixed $z\in\mathbb{R}^{k}$, the system (10) transforms the prior distribution from $t=0$ to the conditional measure $\mu_{z}$ at $t=1$.
+parametrized by a variable $z\in\mathbb{R}^{k}$, such that, for a fixed $z\in\mathbb{R}^{k}$, the system (10) transforms the prior distribution from $t=0$ to the conditional distribution $\mu_{z}$ at $t=1$.
 To achieve this goal, we consider the following objective for learning the vector field in (10):
@@
 $\mathcal{L}^{(N)}(v)=\frac{1}{N}\sum_{n=1}^{N}\left\|v^{(\xi(X^{(n)}))}\big((1-t_{n})X_{0}^{(n)}+t_{n}X^{(n)},t\big)-(X^{(n)}-X^{(n)}_{0})\right\|^{2}\,$, (13)
-where $t_{1},t_{2},\dots,t_{N}$ are drawn independently from $U([0,1])$, $X_{0}^{(1)},X_{0}^{(1)},\dots,X_{0}^{(N)}$ are drawn independently from $\nu_{\mathrm{prior}}$, and $X^{(1)},X^{(2)},\dots,X^{(N)}$ are the data points.
+where $t_{1},t_{2},\dots,t_{N}$ are drawn independently from $U([0,1])$, $X_{0}^{(1)},X_{0}^{(1)},\dots,X_{0}^{(N)}$ are drawn independently from the prior $\nu_{\mathrm{prior}}$, and $X^{(1)},X^{(2)},\dots,X^{(N)}$ are the data points.
 3.2 Enhanced learning approach with biased trajectory data
-We consider the case where the target distribution $\mu$ is the Boltzmann–Gibbs distribution in (1), which is the invariant distribution of a diffusion process $x_{t}$ in $\mathbb{R}^{d}$.
-For simplicity, we assume that $x_{t}$ is governed by the overdamped Langevin dynamics
+We consider the case where the target distribution $\mu$ is the Boltzmann–Gibbs distribution with a density in (1). Let us introduce the overdamped Langevin dynamics
 $dx_{t}=-\nabla V(x_{t})\,dt+\sqrt{2\beta^{-1}}\,dw_{t}\,,\quad t\geq 0\,$, (14)
 where $V:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is a smooth potential function that increases to infinity sufficiently fast as $|x|\rightarrow+\infty$ and $w_{t}$ is a standard Brownian motion in $\mathbb{R}^{d}$.
-In this case, it is known that the Boltzmann–Gibbs distribution $\mu$ in (1) is invariant under the process $x_{t}$ in (14) and that $x_{t}$ is ergodic with respect to $\mu$.
-Thanks to the ergodicity of the process, the empirical objective can be evaluated from trajectory data $X^{(1)},X^{(2)},\dots,X^{(N)}$ obtained by sampling (14) using a numerical scheme at time $t=h,2h,\dots,Nh$, where $h$ denotes the step-size and $N$ is the total number of sampling steps.
-We are particularly interested in the case where the potential $V$ has multiple local minima and $\beta$ is (relatively) large.
-In this case, the probability density of the invariant distribution $\mu$ is highly non-uniform in space, since the probability that $x_{t}$ is within the vicinity of one of the local minima (metastable regions) is much larger than the probability that $x_{t}$ is elsewhere.
-Likewise, for a proper CV map $\xi$, the pushforward distribution $\widetilde{\mu}$ of $\mu$ defined in (4) is highly non-uniform in $\mathbb{R}^{k}$ as well.
-In this case, the generative model that is learned by minimizing the objective (11) (or its empirical counterpart (13)) using training data sampled from (14) may be inferior for values $z$ that belong to low-density regions under $\widetilde{\mu}$, due to the relatively small contribution of the loss in such regions to the aggregated objective (11) (see (12)).
-In order to overcome the aforementioned issue, we utilize the observation (as can be seen from the derivation (12)) that different probability distributions of $X_{1}$ can be chosen in (11), as long as the corresponding conditional probability distributions coincide with the conditional probability distributions $\mu_{z}$ of $\mu$.
-In particular, let us consider the reweighted (RW) probability distribution defined by
+It is known that $\mu$ is invariant under the process $x_{t}$ in (14) and that $x_{t}$ is ergodic with respect to $\mu$.
+Thanks to the ergodicity of the process, the empirical objective (13) can be evaluated using trajectory data $X^{(1)},X^{(2)},\dots,X^{(N)}$ obtained by sampling (14) with a numerical scheme at time $t=h,2h,\dots,Nh$, where $h$ denotes the step-size and $N$ is the total number of sampling steps.
+We are particularly interested in the case where the potential $V$ has multiple local minima and $\beta$ is relatively large (i.e.
+the temperature is low).
+In this case, the probability density of the invariant distribution $\mu$ is highly non-uniform in space, since the probability that $x_{t}$ is within the vicinity of one of the local minima (metastable regions) is much larger than the probability that $x_{t}$ is found elsewhere.
+Likewise, for a proper CV map $\xi$, the pushforward distribution $\widetilde{\mu}$ defined in (4) is highly non-uniform in $\mathbb{R}^{k}$ as well.
+The generative model that is learned by minimizing the objective (11) (or its empirical counterpart (13)) using training data sampled from (14) may be inferior in the low-density regions under $\widetilde{\mu}$, because of the relatively small contribution of the loss in such regions to the aggregated objective (11) (see (12)).
+In order to overcome the aforementioned issue, we utilize the observation (as can be seen from the derivation (12)) that different probability distributions of $X_{1}$ can be chosen in (11), as long as the corresponding conditional probability distributions coincide with the conditional probability distributions $\mu_{z}$.
+In particular, let us consider the reweighted (RW) probability distribution
 $\mu_{\mathrm{RW}}(dx)=\bar{Z}^{-1}Z^{-1}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx\,,\quad x\in\mathbb{R}^{d}\,$, (15)
 where $A:\mathbb{R}^{k}\rightarrow\mathbb{R}$ and $\bar{Z}=Z^{-1}\int_{\mathbb{R}^{d}}\mathrm{e}^{-\beta\big(V(x)-A(\xi(x))\big)}\,dx$ is a normalizing factor.
-Then, for each $z\in\mathbb{R}^{k}$, it is straightforward to verify that the conditional probability distribution of $\mu_{\mathrm{RW}}$ on the level-set $\Sigma_{z}$ is $\mu_{z}$ given in (3).
+Then, for each $z\in\mathbb{R}^{k}$, it is straightforward to verify that the conditional probability distribution of $\mu_{\mathrm{RW}}$ on the level-set $\Sigma_{z}$ coincides with $\mu_{z}$ given in (3).
 Moreover, using (5) and (6), we can obtain that the pushforward of $\mu_{\mathrm{RW}}$ by the map $\xi$ satisfies
 $\widetilde{\mu}_{\mathrm{RW}}(dz)=\bar{Z}^{-1}\mathrm{e}^{-\beta\big(F(z)-A(z)\big)}dz\,,\quad z\in\mathbb{R}^{k}\,$, (16)
-where $F$ is the free energy (6) associated to the map $\xi$. With the choice of the probability distribution $\mu_{\mathrm{RW}}$ in (15), the objective (11) becomes
+where $F$ is the free energy (6) associated to the map $\xi$. With the choice of the probability distribution $\mu_{\mathrm{RW}}$ in (15), the objective (11) changes to
 $\mathcal{L}_{\mathrm{RW}}(v)=\mathbb{E}_{t\sim U([0,1])}\mathbb{E}_{X_{0}\sim\nu_{\mathrm{prior}},X_{1}\sim\mu_{\mathrm{RW}}}\left\|v^{(\xi(X_{1}))}\big((1-t)X_{0}+tX_{1},t\big)-(X_{1}-X_{0})\right\|^{2}\,$, (17)
-and by analogy with (18) we can derive that
+and similar to (12) we can derive that
 $\mathcal{L}_{\mathrm{RW}}(v)=$ $\mathbb{E}_{z\sim\widetilde{\mu}_{\mathrm{RW}}}\Big(\mathcal{L}_{\mathrm{FM},\mu_{z}}(v^{(z)})\Big)\,.$ (18)
@@
 where $f_{t}(z)$ is an estimation of the gradient of the free energy $F$ (i.e. the mean force), based on the expression [25, Lemma 3.9]:
-$\frac{\partial F}{\partial z_{i}}(z)=\mathbb{E}_{\mu_{z}}\left[\sum_{l=1}^{d}\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\frac{\partial V}{\partial x_{l}}-\beta^{-1}\sum_{l=1}^{d}\frac{\partial}{\partial x_{l}}\bigg(\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\bigg)\right]\,,\quad i=1,2,\dots,k\,$, (21)
+$\frac{\partial F}{\partial z_{i}}(z)=\mathbb{E}_{\mu_{z}}\left[\sum_{l=1}^{d}\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\frac{\partial V}{\partial x_{l}}-\beta^{-1}\sum_{l=1}^{d}\frac{\partial}{\partial x_{l}}\bigg(\sum_{j=1}^{k}(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}\frac{\partial\xi_{j}}{\partial x_{l}}\bigg)\right]\,,~i=1,2,\dots,k\,$, (21)
 where $(\nabla\xi^{\top}\nabla\xi)^{-1}_{ij}$ denotes the $ij$-component of the inverse of the matrix $\nabla\xi^{\top}\nabla\xi\in\mathbb{R}^{k\times k}$.
-In practice, a grid in $\mathbb{R}^{k}$ is defined and the mean value (21) is estimated for each cell of the grid.
-In each simulation step, the estimation is updated for the cell which the current state $\xi(x_{t})$ falls into.
-When the cell has been visited for more than a certain (pre-defined) number of times, the current value of the estimation is used as the biasing term $f_{t}(\xi(x_{t}))$ in (20).
-As $t\rightarrow+\infty$, we can expect that the empirical estimations converge to the true means in (21) and, accordingly, the process (20) converges to its equilibrium counterpart (19) with $A=F$.
+In practice, the mean value (21) is estimated on a grid in $\mathbb{R}^{k}$.
+In each simulation step, the estimation is updated for the cell where the current state $\xi(x_{t})$ belongs to.
+After the cell has been visited for a certain (pre-defined) number of times, the current value of the estimation in that cell is used as the biasing term $f_{t}(\xi(x_{t}))$ in (20).
+As $t\rightarrow+\infty$, the empirical estimations are expected to converge to the true means in (21) and, accordingly, the process (20) is expected to converge to its equilibrium counterpart (19) with $A=F$.
 3.3 Projection step
-Once the model is learned using the approach described in the previous subsection, it can be used to generate new data points on different target level-sets of $\xi$ by integrating the ODE (10) with different values $z\in\mathbb{R}^{k}$.
+Once the model is learned using the approach described in the previous section, it can be used to generate new data points on different level-sets of $\xi$ by integrating the ODE (10) with different values $z\in\mathbb{R}^{k}$.
 However, due to numerical errors, the generated data may not lie exactly on the target level-set.
-Here, we propose to implement an additional post-processing step optionally in order to project the generated data onto the target level-set.
+Here, we propose a post-processing step that can be used optionally (depending on the specific applications) to project the generated data onto the target level-set.
 Also refer to [33] for discussions on projection onto manifolds in the context of generative modeling.
+In the following, we discuss two different projection methods.
 The first method is the projection to the closest point on the level-set, which amounts to solving
-$\bar{x}:=\arg\min_{y\in\Sigma_{z}}\|x-y\|^{2}$ (22)
+$\bar{x}:=\operatorname*{arg\,min}_{y\in\Sigma_{z}}\|x-y\|^{2}$ (22)
 for each generated point $x$.
-However, unless the level-set $\Sigma_{z}$ can be parametrized by certain coordinates, solving (22) is computationally infeasible because it is a constrained (high-dimensional) optimization task.
-Another simpler method is to project the generated state $x$ along the ODE flow [49]
+However, unless the level-set $\Sigma_{z}$ can be parametrized by certain coordinates, solving (22) is computationally infeasible because it involves constraints.
+The second method is to project the generated state $x$ along the ODE flow [49]
 $\frac{dY_{t}}{dt}=-\nabla G(Y_{t})\,,\quad Y_{0}=x\,$, (23)
-where the function $G:\mathbb{R}^{d}\rightarrow\mathbb{R}$ is defined as $G(y):=\frac{1}{2}\|\xi(y)-z\|^{2}$ for $y\in\mathbb{R}^{d}$.
-Thanks to the gradient structure, it can be shown that, under certain regularity assumptions on $\xi$, the quantity $G(Y_{t})$ decreases to zero exponentially fast as $t\rightarrow+\infty$, implying that the state evolved by the flow of (23) approaches the level-set $\Sigma_{z}$.
-In practice, to perform the projection, it suffices to integrate the ODE (23) up to a time when certain criteria is met, i.e.
+where $G(y):=\frac{1}{2}\|\xi(y)-z\|^{2}$ for $y\in\mathbb{R}^{d}$.
+Under certain regularity assumptions on $\xi$, the quantity $G(Y_{t})$, as $t\rightarrow+\infty$, decreases to zero exponentially fast thanks to the gradient structure of (23), implying that the state evolved by the flow of (23) approaches the level-set $\Sigma_{z}$.
+In practice, to perform the projection, it suffices to integrate the ODE (23) up to a time when certain criteria is met, e.g.
 $\|\xi(Y_{t})-z\|<\epsilon_{\mathrm{tol}}$, where $\epsilon_{\mathrm{tol}}>0$ is a small parameter specifying the tolerance.
+It is worth pointing out that, while this projection step is helpful to meet the condition that the states lie on the target level-set, it does not necessarily improve the quality of the generated data, as measured by the distance between the empirical distribution of the data and the target conditional probability distribution on the level-set.
+For systems in physics, it may also produce projected states that are not physically meaningful.
+To avoid such issues, it is necessary that the generative model is well trained, so that the generated data (before projection) is close to the target level-sets.
+Besides, incorporating geometrical or physical information in the projection step (e.g.
+by requiring that the distance between the projected state and the starting state to be small or the energy to be low) could be helpful.
+Finally, instead of a projection step, one could also relax the data points generated on the target submanifold with some constrained Langevin dynamics in order to improve their physical relevance [26, 27].
 4 Experiments
-In this section, we demonstrate the capability of our approach through different representative datasets.
-The section is organized as follows: we first consider a synthetic dataset in $\mathbb{R}^{2}$ to illustrate the method.
-Next, we evaluate its performance on the trajectory data of the system under the Müller–Brown potential, a widely used benchmark system exhibiting metastability.
-Finally, we apply the approach to the alanine dipeptide molecular system, showcasing its applicability to molecular systems with complex energy landscapes.
-The neural network and the training were implemented using PyTorch [39].
+In this section, we demonstrate the capability of our method by applying it to different representative datasets.
+First, we consider a synthetic dataset in $\mathbb{R}^{2}$ to illustrate the method.
+Next, we evaluate the performance of our method on the trajectory data of the process under the Müller–Brown potential, which is a widely used benchmark system exhibiting metastability.
+Finally, we apply our method to the alanine dipeptide molecular system.
+The neural networks and the training were implemented using PyTorch [39].
 4.1 2D Synthetic Dataset
-Figure 1: (a) Scatter plot of the original 2D dataset sampled using the make_circles function from scikit-learn [40]. (b) Density estimate of the corresponding CV map $\xi(x)=x_{1}^{2}+x_{2}^{2}$.
+Figure 1: (a) Scatter plot of the 2D dataset sampled using the make_circles function from scikit-learn [40]. (b) Density estimate of the corresponding CV map $\xi(x)=x_{1}^{2}+x_{2}^{2}$.
 Figure 2: Results for the 2D dataset.
-(a) Mean value of $\xi$ of generated samples compared to the intended CV value.
+(a) Mean value of $\xi$ on the generated samples compared to the intended CV value.
 (b) Mean deviation of $\xi$ on generated samples from the target CV value.
-(c) Proportion of samples with positive $x_{1}$ values in the original and generated samples for different target CV values.
-Figure 3: Evolution of the two-dimensional samples generated by the conditional ODE flow for a fixed target CV value $z=0.6$.
-Each panel shows the particle positions at different integration times $t$, illustrating how the initial Gaussian cloud progressively morphs into the target distribution.
-The 2D dataset (Figure 1(a)) is sampled using the make_circles function from the package scikit-learn [40].
+(c) Proportion of samples with positive $x_{1}$ values in the original and in the generated samples for different target CV values.
+Figure 3: Evolution of the two-dimensional samples under the ODE flow for a fixed target CV value $z=0.6$.
+The panels show the particle positions at different integration times $t$, illustrating how the initial Gaussian cloud progressively morphs into the target distribution.
+We study the 2D dataset (Figure 1(a)) that is sampled using the make_circles function from the package scikit-learn [40].
 The CV map is defined as $\xi(x)=x_{1}^{2}+x_{2}^{2}$ for $x=(x_{1},x_{2})\in\mathbb{R}^{2}$, whose empirical density is shown in Figure 1(b).
@@
 The vector field $v$ in (10) is parameterized as a fully connected feedforward neural network, comprising a total of four layers: an input layer, two hidden layers with $128$ neurons each and Tanh activations, followed by an output layer of size $2$.
-The input consists of the spatial coordinate $x$ in $\mathbb{R}^{2}$, a time scalar $t$, and the target CV value $z$ (with dimension 1).
-To learn the vector field $v$, the empirical objective $\mathcal{L}^{(N)}(v)$ in (13) is minimized using the Adam optimizer for up to 1000 epochs with early stopping (patience of 200), a learning rate of $10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $1000$.
+The input consists of the spatial coordinate $x$ in $\mathbb{R}^{2}$, a time scalar $t$, and the target CV value $z$ (of dimension 1).
+To learn the vector field $v$, the empirical objective $\mathcal{L}^{(N)}(v)$ in (13) was minimized using the Adam optimizer for up to 1000 epochs with early stopping (patience of 200 epochs), a learning rate of $10^{-3}$, a weight decay of $10^{-4}$, and a batch size of $1000$.
 For other parameters, their default values in PyTorch were used.
-Next, we evaluate the performance of the learned model for generating states conditioned on the value $z$.
+Next, we evaluate the performance of the learned model for generating states conditioned on the value $z$ of the map $\xi$.
 To this end, we choose the values of $z$ uniformly within an extended range around the values visited in the dataset (with a 5% margin), and for each chosen value of $z$, $1000$ new states are generated by integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $1000$ time steps, starting from the standard Gaussian prior at $t=0$.
@@
 where $(X^{(i)})_{1\leq i\leq N}$ are the generated samples, as well as the proportion of states with $x_{1}\geq 0$, which are shown in Figures 2(a)2(b), and 2(c), respectively.
-The results in Figure 2 demonstrate that the proposed model performs consistently well across all evaluation metrics in the two-dimensional case.
-The generated CV values closely match the intended targets, the deviation remains low across the entire range of values, and the model accurately learns the conditional distribution.
-Finally, Figure 3 presents the temporal evolution of the generated distribution for $z=0.6$, obtained by integrating the ODE from a standard Gaussian initial cloud of $2000$ points.
+The results in Figure 2 demonstrate that the trained model performs consistently well across all evaluation metrics.
+Namely, the generated CV values closely match the intended targets, the deviation remains low across the entire range of values, and the model accurately learns the conditional distribution.
+Finally, Figure 3 presents the temporal evolution of the generated distribution for $z=0.6$, obtained by integrating the ODE from an initial cloud of $2000$ points sampled from the standard Gaussian prior.
 Successive snapshots are taken at increasing integration times $t\in\{0,0.25,0.50,0.75,1.0\}$, showing how the initially isotropic Gaussian distribution gradually transforms into the target ring-shaped configuration.
@@
 low-potential regions), separated by a shallow well.
-At low temperatures, the corresponding dynamics exhibits metastability with a curved transition pathway, making it an ideal model system for studying the complex behavior of metastable dynamics.
+At low temperatures, the corresponding process (14) exhibits metastability with a curved transition pathway, making it an ideal model system for studying the complex behavior of metastable dynamics.
 Model Description and Training.
-The dataset is prepared by sampling the trajectory of the overdamped Langevin dynamics (14) up to time $t=600$ (corresponding to $3\times 10^{6}$ steps), with the integration time step $\Delta t=2\times 10^{-4}$ and inverse temperature $\beta=0.1$.
-The starting point is chosen in the deepest well, with configuration (0.6,1.2)(-0.6,1.2).
-By recording states every 100 steps, we obtained a dataset consisting of $3\times 10^{4}$ configurations.
+The dataset was prepared by sampling the trajectory of the overdamped Langevin dynamics (14) up to time $t=600$ (corresponding to $3\times 10^{6}$ steps), with the integration time step $\Delta t=2\times 10^{-4}$ and the inverse temperature $\beta=0.1$, starting from the state (0.6,1.2)(-0.6,1.2) in the deepest well.
+A dataset consisting of $3\times 10^{4}$ states was obtained by recording the states every 100 steps.
 Figure 4: Heat map and contour representation of the Müller–Brown potential landscape.
-The two major low-potential regions and the shallow low-potential region are shown in dark blue and light blue, respectively, while the contour lines highlight specific energy levels.
+The two major low-potential regions and the shallow low-potential region are shown in dark blue and light blue, respectively, while the contour lines highlight the specific potential energy levels.
 In order to define a CV map, we first trained an autoencoder with a one-dimensional bottleneck, modeled as a fully connected feedforward neural network, by minimizing the mean squared reconstruction error on the generated trajectory data.
@@
 The CV map $\xi:\mathbb{R}^{2}\rightarrow\mathbb{R}$ is defined by the trained encoder.
-Figure 5(a) shows the level-sets of $\xi$ with the generated trajectory data superimposed, whereas Figure 5(b) shows the scatter plot of the potential energy $V(x)$ for each state $x$ in the trajectory against the encoded CV value $\xi(x)$.
-To test the enhanced learning technique discussed in Section 3.2, we also employed the ABF dynamics (20) with the learned CV map $\xi$, where the biasing force $f_{t}$ was computed based on (21) on a uniform grid with grid spacing $0.1$ that covers the range [2.6,3.5][-2.6,3.5].
-To ensure accurate estimation, the biasing force in a cell of the grid is applied to the system only after 100 samples have been used to estimate (21).
+Figure 5(a) shows the level-sets of $\xi$ with the sampled trajectory data superimposed, whereas Figure 5(b) shows the scatter plot of the potential energy $V(x)$ for each state $x$ in the trajectory against the encoded CV value $\xi(x)$.
+To test the enhanced learning technique discussed in Section 3.2, we also employed the ABF dynamics (20) with the learned CV map $\xi$, where the biasing force $f_{t}$ was estimated according to (21) on a uniform grid with grid spacing $0.1$ that covers the range [2.6,3.5][-2.6,3.5].
+To ensure accurate estimation, the biasing force in a grid cell was applied to the system after the cell has been visited 100 times (equivalently, 100 samples have been collected to estimate (21)).
 The biased simulation was run for $1.5\times 10^{5}$ steps and, after the initial $5\times 10^{4}$ simulation steps (equilibration), states were recorded every $10$ steps, resulting in a trajectory dataset of size $10^{4}$.
-Similarly to Figures 5(a)5(b), Figure 6(a) and 6(b) show the scatter plot of the biased trajectory data superimposed on the level-sets of $\xi$ and the scatter plot of potential values $V(x)$ against the encoded CV value $\xi(x)$, respectively.
-We trained our generative model (10) using the generated trajectory data and the learned CV map $\xi$.
+Similar to Figures 5(a)5(b), Figures 6(a) and 6(b) show the scatter plot of the ABF-biased trajectory data superimposed on the level-sets of $\xi$ and the scatter plot of potential values $V(x)$ against the encoded CV value $\xi(x)$, respectively.
+We trained the ODE model (10) using the generated trajectory data and the learned CV map $\xi$.
 The vector field $v$ is modeled by a fully connected feedforward neural network, which consists of an input layer of size 4 (i.e.
 the total dimension of the state $x\in\mathbb{R}^{2}$, time $t\in[0,1]$ and the target CV value $z\in\mathbb{R}$), three hidden layers of size 128 with Tanh activation, and an output layer of size $2$.
-Training was performed for 3000 epochs by minimizing the objective (13) (respectively, the empirical counterpart of (17) for biased trajectory data) using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $512$.
+Training was performed for 3000 epochs by minimizing the objective (13) using the Adam optimizer with a learning rate of $10^{-3}$ and a batch size of $512$.
 Results.
-We evaluated the trained generative models by generating new states with prescribed target CV values 2.0,0.0-2.0,0.0 and $2.0$.
+We evaluated the trained generative models by generating new states with prescribed target CV values $z\in\{-2.0,0.0,2.0\}$.
 For each target CV value $z$, $10^{3}$ states sampled from the two-dimensional standard Gaussian distribution were propagated by numerically integrating the ODE (10) using the torchdiffeq library [3] (4th-order Runge-Kutta scheme) up to $t=1$ with $100$ uniform time steps.
@@
 These samples are visualized in Figures 7(a) and 8(a).
-Comparing these two figures reveals that the model trained using biased trajectory data produces spatially consistent samples distributed along the correct CV contours (Figure 8(a)), whereas the model trained using unbiased trajectory data tends to generate inconsistent samples when the target level-sets are not well covered by the training data (i.e.
+Comparing these two figures reveals that the model trained using the ABF-biased trajectory data produces spatially consistent samples distributed along the correct CV contours (Figure 8(a)), whereas the model trained using the unbiased trajectory data tends to generate inconsistent samples when the target level-sets are not well covered by the training data (i.e.
 level-sets corresponding to z=2.0z=-2.0 and $z=0.0$ in Figure 7(a)).
-To quantify the evaluation, the encoded CVs of generated samples were compared to the target value $z$, and the root mean squared deviation defined in (24) was computed for a uniform grid of target values $z\in[-2.9,3.6]$.
-The deviation curves in Figure 9 summarize this evaluation: the model trained on biased trajectories achieves a lower deviation across the CV range.
-Figure 10 shows snapshots of the temporal evolution of the point distributions under the learned flow model (10), taken at specific time points $t=0.0,0.5,0.9,0.95,1.0$, illustrating the transformation from the initial standard Gaussian distribution to the final distribution conditioned on the latent variable $z$.
-Figure 5: (a) Trajectory samples overlaid on level-sets of the learned CV map $\xi$.
+To quantify the evaluation, the CV values of the generated samples were compared to the target value $z$, and the root mean squared deviation defined in (24) was computed for the target values $z$ on a uniform grid of the interval [2.9,3.6][-2.9,3.6].
+The deviation curves in Figure 9 summarize this evaluation: the model trained on the ABF-biased trajectory achieves a lower deviation across the CV range.
+Figure 10 shows snapshots of the temporal evolution of the point distributions under the learned ODE model (10), taken at specific time points $t=0.0,0.5,0.9,0.95,1.0$, illustrating the transformation from the initial standard Gaussian distribution to the final distribution conditioned on the latent variable $z$.
+Figure 5: (a) Trajectory samples from unbiased sampling overlaid on level-sets of the learned CV map $\xi$.
 The color bar indicates the corresponding CV values across the configuration space.
 (b) Scatter plot of the potential energy versus the CV value.
-Figure 6: (a) Biased trajectory samples obtained from ABF simulation overlaid on level-sets of the learned CV map $\xi$.
+Figure 6: (a) Trajectory samples obtained from the ABF simulation overlaid on level-sets of the learned CV map $\xi$.
 The color bar indicates the corresponding CV values across the configuration space.
-(b) Scatter plot of the potential energy versus the CV value for states in the biased trajectory.
-Figure 7: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on unbiased trajectory data. (a) Without projection. (b) With projection.
-Figure 8: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on biased trajectory data. (a) Without projection. (b) With projection.
-Figure 9: Comparison of the deviations defined in (24) for different target CV values between the model trained on unbiased trajectory data and the model trained on biased trajectory data (ABF).
+(b) Scatter plot of the potential energy versus the CV value for states in the ABF-biased trajectory.
+Figure 7: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on the unbiased trajectory data. (a) Without projection. (b) With projection.
+Figure 8: Generated samples for different CV values $z\in\{-2.0,0.0,2.0\}$ using the model trained on the ABF-biased trajectory data. (a) Without projection. (b) With projection.
+Figure 9: The deviation defined in (24) as a function of the target CV value for the model trained on the unbiased trajectory data and the model trained on the ABF-biased trajectory data.
 We also tested the method introduced in Section 3.3 for projecting the generated samples.
-For each target CV value $z$, the samples obtained from the generative models (as shown in Figures 7(a) and 8(a)) were propagated by numerically integrating the ODE 23 using the explicit Euler scheme, where we used step size $0.01$, the maximal number of integration steps $7000$ and tolerance $\epsilon_{\mathrm{tol}}=10^{-3}$.
+For each target CV value $z$, the samples obtained from the generative models (as shown in Figures 7(a) and 8(a)) were propagated by integrating the ODE (23) numerically using the explicit Euler scheme, where we used step size $0.01$, the maximal number of integration steps $7000$ and tolerance $\epsilon_{\mathrm{tol}}=10^{-3}$.
 The resulting states are shown in Figures 7(b) and 8(b), respectively, from which we can conclude that the generated samples were successfully projected onto the corresponding level-sets.
@@
 We chose z=2.0z=-2.0 and sampled $5\times 10^{4}$ states on the corresponding level-set using the constrained sampling scheme in [5] with a step size $2.0\times 10^{-4}$.
-Each state $x$ was assigned the weight $|\nabla\xi(x)|^{-1}$, which is proportional to the likelihood ratio between the conditional distribution (3) and the invariant distribution of the constrained numerical scheme in the continuous-time limit, so that the weighted samples can be used to approximate the true conditional distribution on the level-set (see [5, 49] for detailed discussions).
-With these data, we estimated the densities of the first component $x_{1}$ for the samples generated by the learned generative model, their projections on the level-set by the method in Section 3.3, and the weighted samples obtained by constrained sampling.
-The results are shown in Figure 11(a) and Figure 11(b) for the model trained using unbiased trajectory data and the model trained using biased trajectory data, respectively.
-From these two figures, we can conclude that the projection step maps the generated states onto the level-set without modifying their distribution.
-Moreover, it can be observed that, compared to the model learned using unbiased trajectory data, the generative model learned using biased trajectory generates data whose distribution is closer to the true conditional distribution.
-We also computed probability densities of other quantities such as the potential energy and distance to the mean.
+In order to approximate the true conditional distribution, each of the sampled states $x$ was assigned the weight $|\nabla\xi(x)|^{-1}$, which is proportional to the likelihood ratio between the conditional distribution (3) and the invariant distribution of the constrained numerical scheme in the continuous-time limit (see [5, 49] for detailed discussions).
+With these data, we estimated the densities of the first component $x_{1}$, using the samples generated by the learned generative model, their projections on the level-set by the method in Section 3.3, and the weighted samples obtained by constrained sampling.
+The results are shown in Figure 11(a) and Figure 11(b) for the model trained using the unbiased trajectory data and the model trained using the ABF-biased trajectory data, respectively.
+From these two figures, we can conclude that the projection maps the generated states onto the level-set without modifying their distribution significantly.
+Moreover, it can be observed that, compared to the model trained using the unbiased trajectory data, the ODE model trained using the ABF-biased trajectory generates data whose distribution is closer to the true conditional distribution.
+We also computed probability densities of other quantities such as the potential energy and the distance to the mean.
 The results are similar to those in Figure 11 and therefore are not shown here.
 Figure 10: Time evolution of the data points under the learned model (10) for three different target values of $z\in\{-2.0,0.0,2.0\}$.
-The top part of the figure corresponds to the model trained with unbiased trajectory data, and the bottom part corresponds to the model trained with biased trajectory data.
+The top part of the figure corresponds to the model trained with the unbiased trajectory data, and the bottom part corresponds to the model trained with the ABF-biased trajectory data.
 Contours and filled colormaps show the level-sets and the profile of the CV map $\xi$, respectively.
 Points corresponding to each $z$ value are drawn in distinct colors to track their evolution over time.
-Figure 11: Density estimations of the first component $x_{1}$ for the samples generated using the generative model for the value z=2.0z=-2.0 (without projection), their projections onto the level-sets (with projection), and the samples obtained by constrained sampling on the level-set (reference).
-(a) Results for the generative model trained using unbiased trajectory data.
-(b) Results for the generative model trained using biased trajectory data.
+Figure 11: Density estimations of the first component $x_{1}$ for the samples generated using the generative model for the target CV value z=2.0z=-2.0 (without projection), their projections onto the level-sets (with projection), and the samples obtained by constrained sampling on the level-set (reference).
+(a) Results for the generative model trained using the unbiased trajectory data.
+(b) Results for the generative model trained using the ABF-biased trajectory data.
 4.3 Alanine Dipeptide
 Model Description and Training.
-In this study, we applied our generative approach to alanine dipeptide, which is a small benchmark molecular system consisting of 22 atoms.
-Our goal is to train the ODE model (10) for generative modeling of configurations of the system’s 10 non-Hydrogen atoms conditioned on its two backbone dihedral angles $\phi$ and $\psi$.
-Trajectory data of the alanine dipeptide system were generated by performing molecular dynamics simulations for $t=1500\,\textrm{ns}$ using the GROMACS package [47], where we adopted the same simulation parameters as in the previous work [24] and the states were stored every $10$ ps, resulting in a trajectory dataset of $1.5\times 10^{5}$ states.
-Only Cartesian coordinates of the 10 non-Hydrogen atoms were then extracted using MDAnalysis [12, 37] and aligned to a reference configuration via the Kabsch algorithm [20], thereby eliminating global translational and rotational motions.
+We applied our generative approach to alanine dipeptide, which is a small benchmark molecular system consisting of 22 atoms.
+We studied two different generative modeling tasks.
+In the first task, our goal is to train the ODE model (10) to generate configurations of the system’s 10 non-Hydrogen atoms conditioned on its two backbone dihedral angles $\phi$ and $\psi$.
+In the second task, our goal is to train the ODE model (10) to generate configurations of the system’s 10 non-Hydrogen atoms conditioned on the dihedral angle $\phi$ alone.
+Trajectory data of the alanine dipeptide system in water was generated by performing molecular dynamics simulations for $t=1500\,\textrm{ns}$ using the GROMACS package [47], where we adopted the same simulation parameters as in the previous work [24] and the states were recorded every $10$ ps, resulting in a trajectory dataset of $1.5\times 10^{5}$ states.
+Cartesian coordinates of the 10 non-Hydrogen atoms were then extracted using MDAnalysis [12, 37] and aligned to a reference configuration via the Kabsch algorithm [20], thereby eliminating global translational and rotational motions.
 The aligned coordinates were then standardized by subtracting the empirical mean and dividing by the empirical standard deviation before being used as training data.
 Figure 12(a) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles.
-In addition to the unbiased MD data, we also generated an independent trajectory by performing an ABF simulation for $t=500\,\textrm{ns}$ along the backbone dihedral angles $\phi$ and $\psi$.
-The ABF method applies a time-dependent biasing force to enhance the sampling of rarely visited regions of conformational space.
-The simulation was performed using the ColVars module [11] and the GROMACS package.
-The states were stored every $10$ steps and the resulting trajectory dataset consists of $5\times 10^{4}$ states.
-Figure 12(b) shows the empirical density of the trajectory data projected onto the space of the two dihedral angles.
-All ABF-biased coordinates were preprocessed in the same manner as the unbiased data—centered, aligned to a reference structure, flattened, and then standardized.
-For the modeling of the vector field $v$ in (10), we implemented a fully connected feedforward neural network that takes as input the atomic coordinates $x\in\mathbb{R}^{30}$, the temporal parameter $t\in[0,1]$, and the CV value $z\in\mathbb{R}^{2}$, which correspond to the standardized values of $(\phi,\psi)$.
-The feedforward network consists of five hidden layers with 512 hidden units and Tanh activations, while the output layer returns the vector field in $\mathbb{R}^{30}$.
-The training of the vector field $v$ was performed by minimizing the objective (13) for 1000 epochs, using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of 512, and default PyTorch parameters (e.g., for momentum coefficients, numerical stability parameter, and weight decay).
-After training, new samples in $\mathbb{R}^{30}$ can be generated conditioned on the value of $z$ (i.e.
-the target value of $(\phi,\psi)$) by integrating the ODE (10) from random initial positions sampled from the standard Gaussian distribution.
+In addition to the unbiased trajectory, we also generated two biased trajectory datasets of the system by performing ABF simulations using the ColVars module [11] and the GROMACS package.
+Specifically, for the first task (i.e.
+the generation conditioned on both dihedral angles $\phi$ and $\psi$), we performed an ABF simulation for $t=500\,\textrm{ns}$, where the CV was defined as the two dihedral angles $\phi$ and $\psi$.
+The states were stored every $10$ ps and the resulting trajectory dataset consists of $5\times 10^{4}$ states.
+For the second task (i.e.
+the generation conditioned on the dihedral angle $\phi$ alone), we ran an ABF simulation for $t=100\,\textrm{ns}$, where the CV was defined as the backbone dihedral angle $\phi$ alone.
+The states were stored every $1$ ps, resulting in a trajectory dataset of $10^{5}$ states.
+Figures 12(b) and 12(c) show the empirical densities of the two trajectory datasets projected onto the space of the two dihedral angles, respectively.
+It can be observed that different values of the corresponding CVs are more uniformly sampled under the ABF simulations.
+Figures 13(a) and 13(b) display the free energy profiles associated to the two corresponding CVs (reconstructed using the ColVars module), but this information on free energies is not used in the study below.
+Finally, coordinates of states in both datasets were preprocessed in the same manner as the unbiased data—centered, aligned to a reference structure, flattened, and then standardized.
+Figure 12: Empirical densities of the dihedral angles of the trajectory data.
+(a) Trajectory from the unbiased simulation.
+(b) Trajectory from the ABF simulation, where the CV is defined as the dihedral angles $(\phi,\psi)$.
+(c) Trajectory from the ABF simulation, where the CV is defined as the dihedral angle $\phi$.
+Figure 13: (a) Free energy profile associated to the two dihedral angles $(\phi,\psi)$. (b) Free energy profile associated to the dihedral angle $\phi$.
+For the modeling of the vector field $v$ in (10), we implemented a fully connected feedforward neural network that takes as input the atomic coordinates $x\in\mathbb{R}^{30}$, the temporal parameter $t\in[0,1]$, and a value $z$ that corresponds to the conditioning (either angles or their trigonometric functions).
+Specifically, in the first task, the value $z$ is defined either as the standardized values of $(\phi,\psi)$ in $\mathbb{R}^{2}$, or as $z=(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$.
+In the second task, $z$ is defined either as the standardized value of $\phi$ in $\mathbb{R}$, or as $z=(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$.
+The advantage of using trigonometric functions is that the vector field is guaranteed to be periodic as a function of angles.
+In all cases, the feedforward network consists of five hidden layers with 512 hidden units and Tanh activations, while the output layer returns the vector field in $\mathbb{R}^{30}$.
+The training of the vector field $v$ was performed by minimizing the objective (13) for 1000 epochs, using the Adam optimizer with a learning rate of $10^{-3}$, a batch size of 512, and default PyTorch parameters (e.g.
+momentum coefficients, numerical stability parameter, and weight decay).
+For each of the two tasks, the training was carried out both on the unbiased dataset and on the ABF-biased dataset.
+After training, new samples in $\mathbb{R}^{30}$ can be generated conditioned on the target CV value (i.e.
+the target value of $(\phi,\psi)$ in the first task, or the target value of $\phi$ in the second task) by integrating the ODE (10) from random initial positions sampled from the standard Gaussian distribution.
 The generated samples are then denormalized and reshaped to obtain three-dimensional coordinates of the $10$ non-Hydrogen atoms.
-Results.
-To assess the quality of the generated conformations, we chose the target dihedral angles $(\phi,\psi)\in\{(-75^{\circ},150^{\circ}),(52.49^{\circ},34.95^{\circ}),(57.5^{\circ},-126.78^{\circ})\}$, which correspond respectively to the densely populated region, a secondary metastable region, and an isolated point in a low-population region in the dihedral angle space (see Figure 12(a)).
-For each of these values $(\phi,\psi)$, we generated 1000 samples and identified the (reference) configuration in the trajectory whose dihedral angles are closest to $(\phi,\psi)$.
-Root mean square deviation (RMSD) values of the generated samples from the reference were computed after Kabsch alignment.
-Histograms of RMSD values were constructed with fixed bin widths ($0.02$ Å) to visualize the distribution of deviations in different regions represented by these values of dihedral angles.
-These distributions are shown for the model trained using the unbiased dataset in Figure 13 and for the model trained using the biased dataset in Figure 14, respectively.
-RMSD distributions between nearby real structures are shown for the unbiased trajectory dataset in Figures 15(a) and 15(b), serving as a baseline for intrinsic structural variability.
-In the left panel (Figure 15(a)), all pairs of configurations whose dihedral angles $(\phi_{1},\psi_{1})$ and $(\phi_{2},\psi_{2})$ satisfy $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1^{\circ}$ were identified using a nearest neighbors search in the dihedral angle space.
-For each pair, the RMSD was computed after Kabsch alignment to remove global translations and rotations.
-The right panel (Figure 15(b)) extends this analysis by binning the $\phi$ and $\psi$ angles of configurations into a $50\times 50$ grid and computing the mean RMSD of all conformations within each bin (empty bins were ignored).
-For the ABF-biased dataset, the same evaluation workflow was applied, and the results were nearly identical to those for the unbiased dataset.
-Comparing the histograms in Figure 13 and Figure 15, we can conclude that the mean deviations predicted by the trained model are of the same order as the intrinsic structural variability observed in the unbiased dataset.
-To assess the angular consistency of the generated configurations across the dihedral space, we computed heatmaps of mean deviations (see (24)) per $(\phi,\psi)$-bin for the generative models trained using both the unbiased and the ABF-biased datasets, as shown in Figure 16.
-For this analysis, the $\phi$ and $\psi$ angles in the range $180^{\circ},180^{\circ}]$ were binned into intervals of equal length $20^{\circ}$, and the centers of each bin were used as target CV values.
+Results for the first task.
+To assess the quality of the generated configurations, we chose the target dihedral angles $(\phi,\psi)\in\{(-70^{\circ},150^{\circ}),(52.5^{\circ},35^{\circ})\}$, which correspond to a state in the densely populated region and a state in the secondary metastable region, respectively (see Figure 12(a)).
+For each of these values $(\phi,\psi)$, a configuration in the unbiased trajectory whose dihedral angles are closest to the target value $(\phi,\psi)$ was identified as a reference, and 2000 samples were generated conditioned on the target angles for each of the trained models (i.e.
+models trained on the unbiased dataset and on the ABF-biased dataset, with conditioning $z$ being either angles or trigonometric functions).
+The root mean square deviation (RMSD) values of the generated configurations from the reference (after Kabsch alignment) were computed.
+Figures 14(a)14(b) show the estimated densities of the RMSDs.
+Moreover, we selected the configurations in the dataset whose dihedral values belong to the same cell as the target dihedral values $(\phi,\psi)$, when binned into the uniform grid of $180^{\circ},180^{\circ}]\times[-180^{\circ},180^{\circ}]$ with width $10^{\circ}$.
+The RMSD values of the selected configurations from the reference were computed and the estimated densities are also shown in Figures 14(a)14(b) for comparison.
+It can be observed that the density profiles of RMSDs computed for the generated configurations are similar to the density profiles computed for the configurations selected from the trajectory datasets, indicating that the trained models are able to produce consistent configurations conditioned on the target dihedral angles.
+The fact that the density profiles are unimodal is also consistent with the fact that the configuration of alanine dipeptide is largely determined by its two dihedral angles.
+To assess the angular consistency of the generated configurations across the dihedral angle space, we computed heat maps of mean deviations (see (24)) per $(\phi,\psi)$-bin for the trained models, as shown in Figure 15.
+For this analysis, the $\phi$ and $\psi$ angles in the domain $180^{\circ},180^{\circ}]\times[-180^{\circ},180^{\circ}]$ were binned using a uniform grid of equal width $20^{\circ}$ (in both directions), and the centers of grid cells were used as target CV values.
 For each target CV value, $N=200$ configurations were generated using the trained ODE models.
-The resulting mean deviations are shown in Figure 16(a) for the model trained using the unbiased trajectory data, where small deviations are concentrated in the high-probability regions of the CV space, and in Figure 16(b) for the model trained using the ABF-biased trajectory data, which exhibit uniformly low deviations across the full CV space.
-Similarly to the previous example, the projection method introduced in Section 3.3 can be applied as a post-processing step to modify the generated configurations so that they satisfy the constraints on the target dihedral angles.
-We refer to the previous work [34, Appendices B.2 and B.4] for related numerical study.
-Figure 12: Empirical densities of the dihedral angles of the trajectory data. (a) Trajectory from unbiased simulation. (b) Trajectory from ABF simulation.
-Figure 13: Results of the generative model trained with the unbiased dataset.
-Distributions of RMSD values (in Å) for generated samples at three representative CV locations: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$ in a well-populated region (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$ in the secondary metastable basin, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$ corresponding to an isolated point in the low-population region.
-Figure 14: Results of the generative model trained with the ABF-biased dataset.
-Distributions of RMSD values (in Å) for generated samples at the same representative CV locations as in Figure 13: (a) $(\phi,\psi)=(-75^{\circ},150^{\circ})$, (b) $(\phi,\psi)=(52.49^{\circ},34.95^{\circ})$, and (c) $(\phi,\psi)=(57.5^{\circ},-126.78^{\circ})$.
-Figure 15: Analysis of intrinsic structural variability in the unbiased real dataset.
-(a) Distribution of RMSD values (in Å) between pairs of nearby configurations satisfying $(\phi_{1}-\phi_{2})^{2}+(\psi_{1}-\psi_{2})^{2}<1$.
-(b) Histogram of mean RMSD values computed for each $50\times 50$ $(\phi,\psi)$-bin in the dihedral angle space, capturing spatial variations in local conformational variability.
-Figure 16: Heat maps of the mean deviation between the target CV value $(\phi,\psi)$ and the dihedral angles of generated configurations.
-(a) Unbiased molecular dynamics data.
-(b) ABF-biased data.
-Each pixel represents the mean circular distance in degrees, quantifying the accuracy with which the conditional generative model reproduces the desired angular constraints.
+Figure 15(a) and Figure 15(b) show the resulting mean deviations for the models trained on the unbiased trajectory data, where the standardized dihedral angles and the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ were passed to the neural network, respectively.
+For both models, small deviations are concentrated in the high-probability regions of the CV space.
+Comparing these two figures, it is clear that using trigonometric functions is advantageous, as it leads to lower mean deviations in the region close to the right boundary $\phi=180^{\circ}$ in the CV space, since the data points close to the left boundary $\phi=-180^{\circ}$ can contribute to the learning in this region thanks to the periodicity of the vector field (see Figure 12(a)).
+Similarly, Figures 15(c) and 15(d) show the mean deviations for the models trained on the ABF-biased trajectory data, where the standardized dihedral angles and their trigonometric functions were used as inputs of the neural network, respectively.
+Comparing to the models trained on the unbiased trajectory data, lower deviations are observed across the CV space for the models trained on the ABF-biased trajectory data.
+Since in this case the CV space is well covered by the ABF-biased trajectory data, only slight improvement is achieved when trigonometric functions are used (Figure 15(d)).
+Figure 14: Density estimations of RMSD values (in Å) for two target dihedral angles: (a) $(\phi,\psi)=(-70^{\circ},150^{\circ})$ and (b) $(\phi,\psi)=(52.5^{\circ},35^{\circ})$ in the first task.
+In both figures, the line “unbiased, ref” (resp.
+“biased, ref”) corresponds to the density of RMSDs computed for the configurations in the unbiased (resp.
+ABF-biased) trajectory dataset whose dihedral angles are close to the target dihedral values (see the text for details).
+The line with label “unbiased” (resp.
+“biased”) corresponds to the model trained on the unbiased (resp.
+ABF-biased) dataset, where the standardized angles are directly passed to the neural network.
+The line with label “unbiased, tri” (resp.
+“biased, tri”) corresponds to the model trained on the unbiased (resp.
+ABF-biased) dataset, where the values $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)$ are passed to the neural network.
+Figure 15: Heat maps of the mean deviation (i.e.
+the mean circular distance in degrees) between the target CV value $(\phi,\psi)$ and the dihedral angles of generated configurations in the first task.
+(a) Model trained on the unbiased trajectory data, where the standardized dihedral angles are used as input of the neural network.
+(b) Model trained on the unbiased trajectory data, where the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ are used as input of the neural network.
+(c) Model trained on the ABF-biased trajectory data, where the standardized dihedral angles are used as input of the neural network.
+(d) Model trained on the ABF-biased trajectory data, where the trigonometric functions $(\cos\phi,\sin\phi,\cos\psi,\sin\psi)\in\mathbb{R}^{4}$ are used as input of the neural network.
 Lower deviation indicates higher fidelity of the generated ensemble to the target CVs.
+Results for the second task.
+Regarding the generative modeling task conditioned on the angle $\phi$ alone, we tested the models trained on the unbiased trajectory data (Figure 12(a)) and on the ABF-biased data (Figure 12(c)), and for each case the vector field $v$ takes either the standardized angle $\phi$ or the trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ as inputs for the conditioning.
+The interval $180^{\circ},180^{\circ}]$ was divided into subintervals of equal length $20^{\circ}$ and the centers of the subintervals were chosen as target CV values of $\phi$.
+For each of these target values and for each of the trained models, $200$ samples were generated, based on which the mean deviations (see (24)) were computed.
+The results are shown in Figure 16(a).
+Similar to the results in the first task, when the unbiased trajectory data is used in training, low deviations are achieved for values of $\phi$ that are frequently taken by the configurations in the dataset, whereas the deviations are high for values of $\phi$ that are rarely taken by configurations in the dataset.
+In this case, the use of trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ helps to reduce the deviations for target $\phi$ values that are close to $180^{\circ}$.
+When the ABF-biased trajectory data is used in training, the model accuracy is greatly improved, as the deviations are low across different target values $\phi$.
+In this case, the advantage of using trigonometric functions is less obvious since the entire space of $\phi$ is well covered by the data (Figure 12(c)).
+To further assess the quality of the models, we studied the distribution of $\psi$ conditioned on the target values $\phi=-70^{\circ}$ and $\phi=52.5^{\circ}$.
+For each of these target values and for each of the trained models, 1000 samples were generated, based on which the empirical density of the angle $\psi$ was estimated.
+For comparison, the density of $\psi$ conditioned on the target value $\phi$ was also estimated using configurations in the unbiased trajectory dataset and configurations in the ABF-biased trajectory dataset (by selecting configurations whose angle $\phi$ belongs to the same cell as the target value of $\phi$ when binned using a uniform grid with width $20^{\circ}$).
+The results are shown in Figures 16(b) and 16(c).
+As can be observed there, for each of the trained models, the distribution of the angle $\psi$ of the generated samples matches well with the conditional distributions estimated using configurations in the trajectory datasets.
+Finally, as in the previous example, the projection method introduced in Section 3.3 can be applied as a post-processing step to modify the generated configurations so that they satisfy the constraints on the target dihedral angles exactly.
+We refer to the previous work [34, Appendices B.2 and B.4] for a related numerical study.
+Figure 16: Results for the second task.
+(a) Mean deviation (in degrees) as a function of the target value $\phi$.
+(2) Estimated densities of $\psi$ conditioned on $\phi=-70^{\circ}$.
+(3) Estimated densities of $\psi$ conditioned on $\phi=52.5^{\circ}$.
+In each figure, the line with label “unbiased” (resp.
+“biased”) corresponds to the model trained on the unbiased (resp.
+ABF-biased) trajectory dataset, where the model takes the standardized value of $\phi$ as an input.
+The line with label “unbiased, tri” (resp.
+“biased, tri”) corresponds to the model trained on the unbiased (resp.
+ABF-biased) trajectory dataset, where the model takes the trigonometric functions $(\cos\phi,\sin\phi)\in\mathbb{R}^{2}$ as input.
+In (b) and (c), the line with labels “unbiased, ref” (resp.
+“biased, ref”) corresponds to the density of $\psi$ estimated using configurations in the unbiased (resp.
+ABF-biased) trajectory dataset.
 5 Conclusion
 In this work, we have introduced a general framework for the generative modeling of conditional distributions on level-sets of CVs using flow-matching neural ODEs.
-By conditioning the learned flows on CV values, the model can generate samples across multiple level-sets without requiring explicit geometric information.
-To further enhance accuracy on level-sets in low-probability regions, we proposed to leverage biased trajectory data from enhanced sampling techniques such as ABF.
+By conditioning the learned flows on CV values, the model can generate samples across different level-sets without requiring explicit geometric information.
+To enhance the generation accuracy on level-sets in low-probability regions, we proposed to leverage biased trajectory data from enhanced sampling techniques such as ABF.
 Validation on concrete datasets demonstrates that our approach accurately reproduces conditional distributions and successfully exploits biased data to improve learning in scarcely sampled regions.
-Beyond these results, this study opens several promising directions for future work.
-First, to better assess the effectiveness of the approach, it is meaningful to study the accuracy of the generative model by comparing certain distances between the generated distributions and the conditional distributions on level-sets.
-Second, a natural extension would be to combine the approach proposed in this work with methods that propagate the effective (reduced) dynamics in the CV space [21, 50].
-Such a combination would lead to generative modeling approaches that can generate new trajectories of stochastic processes while taking into account the essential dynamics.
+This study opens several promising directions for future work.
+First, it is interesting to apply the method to more complex molecular systems and study how the method performs (in terms of the amount of required data and the model quality) when the size of the molecular system increases, especially in the case where the target conditional distributions are multimodal.
+For such applications, it could be the case that more advanced architectures (e.g.
+using featurization or transformers) would improve the generative capabilities.
+Second, a natural extension would be to combine the approach proposed in this work with methods for identifying good CV maps and methods for propagating the effective (reduced) dynamics in the CV space [21, 50].
+Such a combination would lead to generative modeling approaches that can generate new trajectories of high-dimensional complex stochastic dynamics while taking the essential dynamics into account.
 Overall, our study provides a general and practical approach for conditional generative modeling on manifolds defined by CVs, offering both methodological flexibility and the potential for applications in molecular modeling and beyond.
 6 Acknowledgements
-Fatima-Zahrae Akhyar acknowledges the Zuse institute Berlin for hosting her internship during which this research was conducted.
-Wei Zhang has been funded through DFG’s “Eigene Stelle” program (project 524086759) and has been supported by the DFG-grant CRC 1114 “Scaling Cascades in Complex Systems” Project Number 235221301.
+Fatima-Zahrae Akhyar acknowledges Zuse institute Berlin for hosting her internship during which this research was conducted.
+Wei Zhang has been funded through DFG’s “Eigene Stelle” program (project 524086759) and has been supported by the Young Investigator Program of the DFG-grant CRC 1114 “Scaling Cascades in Complex Systems” Project Number 235221301.
 Gabriel Stoltz was funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (project EMC2, grant agreement No 810367), and was supported by Hi!